### v2

# Uncertainty-driven Embedding Convolution

Abstract
Text embeddings are essential components in modern NLP pipelines. While numerous embedding models have been proposed, their performance varies across domains. This variability motivates the use of ensemble techniques to combine complementary strengths. However, most existing ensemble methods operate on deterministic embeddings and fail to account for model-specific uncertainty, limiting their robustness and reliability in downstream applications. To address these limitations, we propose Uncertainty-driven Embedding Convolution (UEC). UEC first transforms deterministic embeddings into probabilistic ones in a post-hoc manner. It then computes adaptive ensemble weights based on embedding uncertainty, grounded in a Bayes-optimal solution under a surrogate loss. Additionally, UEC employs an uncertainty-aware similarity function that directly incorporates uncertainty into the similarity scoring, providing a theoretically grounded and efficient surrogate to distributional distances. Extensive experiments on diverse benchmarks demonstrate that UEC consistently improves both performance and robustness by leveraging principled uncertainty modeling. [^1] Code is available at: https://github.com/MLAI-Yonsei/UEC.
1 Introduction
Embeddings are core building blocks in modern NLP, capturing semantic meaning for words, sentences, and documents. They support tasks like similarity (Gao et al., 2021), retrieval (MacAvaney et al., 2019), QA (Devlin et al., 2019), and classification (Cer et al., 2018). Numerous embedding models (Devlin et al., 2019; Liu et al., 2019; Reimers & Gurevych, 2019) have emerged with diverse architectures and training objectives, but their performance varies across tasks and domains. No single model excels universally; instead, models offer complementary strengths. This motivates combining multiple embeddings to leverage their diverse capabilities.
Figure 1: Comparison of deterministic embedding ensemble and UEC. Deterministic averaging (left) uniformly averages embeddings without considering their reliability, often leading to suboptimal decisions. In this example, both candidate embeddings contribute equally, resulting in an incorrect retrieval. In contrast, the proposed UEC (right) adjusts weights based on the uncertainty, assigning higher importance to the more reliable candidate and successfully retrieving the correct answer.
Embedding models can be ensembled at two levels: parameter and representation. At the parameter level, techniques like model merging combine model weights (Wortsman et al., 2022; Ilharco et al., 2022; Yang et al., 2024). However, this approach is often limited by strict architectural constraints. On the other hand, combining embeddings at the representation level, i.e., ensembling the output embeddings themselves, is a more practical and widely applicable strategy. However, most existing ensemble approaches rely on deterministic aggregation methods such as uniform averaging, which fail to consider the reliability or uncertainty of individual embeddings (Khan et al., 2023a; Fang et al., 2019). Without modeling uncertainty, these methods treat all embeddings as equally reliable, which can lead to suboptimal or unstable performance, especially when some models are poorly calibrated or mismatched to the target task. This is evident in Figure 1, where averaging conflicting ‘animal’ and ‘car’ interpretations for ‘jaguar’ leads directly to a retrieval failure.
To address these limitations, we propose Uncertainty-driven Embedding Convolution (UEC), a framework for combining embeddings in a principled, uncertainty-aware manner. UEC consists of three key components. First, it converts pre-trained deterministic embeddings into probabilistic embeddings in a post-hoc fashion, allowing each embedding to represent both its mean and uncertainty. Second, it computes adaptive ensemble weights based on estimated uncertainty, down-weighting less reliable embeddings. This weighting strategy is grounded in a Bayes-optimal solution under a surrogate loss. Third, UEC introduces an uncertainty-aware similarity function that incorporates both distance and variance into the similarity score, offering a theoretically grounded and efficient surrogate to distributional distances. Consequently, empirical results on diverse benchmarks confirm that UEC reliably improves performance and robustness by leveraging principled uncertainty modeling.
In summary, our key contributions are as follows:
  • We propose a post-hoc mechanism to transform pre-trained deterministic embeddings into probabilistic representations, thereby enabling inherent uncertainty quantification for diverse existing embedding models.
  • We introduce the UEC framework, which adaptively combines multiple embeddings by weighting their contributions based on estimated uncertainties, a method theoretically grounded as a Bayes-optimal solution under surrogate loss minimization.
  • We develop an uncertainty-aware similarity function that explicitly incorporates embedding variance into the similarity scoring process, serving as a theoretically grounded and efficient surrogate to distributional distances.
2 Related Works
2.1 Embedding Ensemble
Embedding ensemble methods aim to improve performance and robustness by combining multiple pre-trained models. Shuang et al. (2019) proposed CDWE, which handles polysemy by generating multiple word prototypes through deconvolution. Fang et al. (2019) enhanced knowledge-aware embeddings by re-weighting knowledge-graph edges. Liu et al. (2025) fused BERT (Devlin, 2018) variants using cross-attention guided by pseudo-labels. Sahlgren (2021) distilled multiple encoders into one by matching their averaged outputs. Khan et al. (2023b) explored fusion strategies for domain-specific BERT models, yielding minor improvements. Despite these advances, most approaches rely on deterministic embeddings, limiting their ability to model uncertainty. In contrast, our method ensembles probabilistic embeddings, capturing both distributional information and model diversity through principled uncertainty estimation.
2.2 Probabilistic Embedding
Probabilistic embeddings represent inputs as distributions rather than points, thereby modeling uncertainty in representation learning. Early approaches such as Gaussian (Vilnis & McCallum, 2014), Gaussian mixture (Chen et al., 2015), and Bayesian embeddings (Barkan, 2017) captured word-level distributional semantics. Sen2Pro (Shen et al., 2023) extended this idea to sentences using pre-trained models, but it requires fine-tuning, lacks post-hoc conversion, and does not support scalable similarity. Subsequent works explored multimodal and vision–language tasks: Chun et al. (2021); Chun (2023) improved cross-modal retrieval and image–text alignment, yet remained modality-specific; ProbVLM (Upadhyay et al., 2023) and BayesVLM (Baumann et al., 2024) incorporated uncertainty into frozen vision–language models through adapters or post-hoc strategies, but with limited generality. Other studies investigated hedging against ambiguous inputs (Oh et al., 2018), compositional multimodal retrieval (Neculai et al., 2022), uncertainty-aware multimodal pre-training (Ji et al., 2023), and scalable probabilistic embeddings with Gaussian process latent variable models (Venkataramanan et al., 2025). Our work provides a Gaussian convolution formulation that not only generates probabilistic embeddings in a post-hoc manner but also supports theoretically grounded and lightweight similarity estimation for downstream tasks.
3 Uncertainty-driven Embedding Convolution
We propose Uncertainty-driven Embedding Convolution (UEC), a principled framework for combining multiple embedding models by modeling predictive uncertainty. UEC consists of three stages (Figure 2): (1) post-hoc conversion of deterministic embeddings into probabilistic ones via Laplace approximation (LA) (Section 3.1), (2) Gaussian convolution with uncertainty-aware weights (Section 3.2), and (3) uncertainty-aware similarity scoring (Section 3.3). This enables robust and adaptive ensemble without manual tuning.
Figure 2: Overview of the UEC framework: UEC first transforms deterministic embeddings from multiple encoder models into probabilistic representations using Laplace approximation. These probabilistic embeddings are then adaptively combined by computing uncertainty-driven ensemble coefficients based on per-dimension variances. Finally, similarity is measured using a uncertainty-aware metric that accounts for both the mean and uncertainty of the ensembled embedding.
3.1 Post-hoc Probabilistic Embedding Model
We first transform deterministic embeddings into probabilistic embeddings using LA (MacKay, 1992; Ritter et al., 2018). This allows each embedding to express both mean and uncertainty, which we later exploit for adaptive model combination and similarity estimation.
Laplace Approximation
LA approximates the posterior distribution of model parameters $\mathbf{w}$ as a Gaussian centered at the maximum a posteriori (MAP) estimate. We denote the training dataset as $\mathcal{D}=\{(\mathbf{x},y)\}$, where $\mathbf{x}$ and $y$ are the input and corresponding output, respectively. The posterior is given by Bayes’ theorem:
$p(\mathbf{W}|\mathcal{D})=\frac{p(\mathcal{D}|\mathbf{W})p(\mathbf{W})}{p(\mathcal{D})},$
where $p(\mathcal{D}|\mathbf{W})$ is the likelihood, $p(\mathbf{W})$ is the prior, and $p(\mathcal{D})$ is the evidence. Since this posterior is generally intractable, LA approximates $p(\mathbf{W}|\mathcal{D})$ with a Gaussian centered at $\widehat{\mathbf{W}}$ (the MAP) and precision matrix given by the Hessian $\mathbf{H}_{\widehat{\mathbf{W}}}$:
[…]
$\mathbf{H}_{\widehat{\mathbf{W}}}$ $=\nabla^{2}[-\log p(\mathcal{D}|\mathbf{W})-\log p(\mathbf{W})]\Big|_{\mathbf{W}=\widehat{\mathbf{W}}}$.
This approximates posterior $p(\mathbf{W}|\mathcal{D})\approx\mathcal{N}(\widehat{\mathbf{W}},\mathbf{H}^{-1}_{\widehat{\mathbf{W}}})$.
Gaussian Embedding Generation
Since computing the full Hessian is often impractical for large models, recent work shows that applying diagonal LA to only the last layer yields sufficiently accurate uncertainty estimates with lower computational overhead (Daxberger et al., 2021). Following this line of work, we apply LA with a diagonal covariance approximation only to the final layer of the embedding model. Let $f(\mathbf{x})$ be an embedding model with $L$ layers, and let $\mathbf{W}^{(L)}$ and $\mathbf{h}^{(L-1)}$ denote the parameters and output of the penultimate layer, respectively. In the deterministic setting, the embedding is computed as $\mathbf{z}=\mathbf{W}^{(L)\top}\mathbf{h}^{(L-1)}$.
[…]
$\mathbf{z}(\mathbf{x})\sim\mathcal{N}(\bm{\mu}(\mathbf{x}),\bm{\Sigma}(\mathbf{x}))=\mathcal{N}(\widehat{\mathbf{W}}^{(L)\top}\mathbf{h}^{(L-1)},\mathbf{h}^{(L-1)\top}\mathbf{H}^{-1}_{\widehat{\mathbf{W}}^{(L)}}\mathbf{h}^{(L-1)}).$
3.2 Bayes-optimal Weights for Uncertainty-driven Embedding Convolution
Gaussian Convolution
Suppose we have $K$ independent embedding models, each producing a probabilistic embedding $\mathbf{z}_{k}(\mathbf{x})\sim\mathcal{N}(\bm{\mu}_{k}(\mathbf{x}),\bm{\Sigma}_{k}(\mathbf{x}))$. To integrate these into a single representation, we perform a Gaussian convolution in the embedding space. Formally, we define the convolutional embedding $\mathbf{z}(\mathbf{x})$ as the result of a weighted combination of independent Gaussian variables:
$\mathbf{z}(\mathbf{x})=\sum_{k=1}^{K}\bm{\pi}_{k}(\mathbf{x})\cdot\mathbf{z}_{k}(\mathbf{x}),$
where $\bm{\pi}_{k}(\mathbf{x})$ denotes the convolution coefficient assigned to model $k$, and $\sum_{k=1}^{K}\bm{\pi}_{k}(\mathbf{x})=\mathbf{1}$.
Since each $\mathbf{z}_{k}(\mathbf{x})$ is Gaussian and the coefficients $\bm{\pi}_{k}(\mathbf{x})$ are deterministic, the resulting embedding $\mathbf{z}(\mathbf{x})$ follows a Gaussian distribution as well, according to the properties of linear combinations of independent Gaussians (Shao, 2008):
$\mathbf{z}(\mathbf{x})\sim\mathcal{N}\left(\sum_{k=1}^{K}\bm{\pi}_{k}(\mathbf{x})\bm{\mu}_{k}(\mathbf{x}),\sum_{k=1}^{K}\bm{\pi}_{k}^{2}(\mathbf{x})\bm{\Sigma}_{k}(\mathbf{x})\right).$ (1)
Gaussian convolution aggregates the centers (means) and spreads (covariances) of the input Gaussians into a single distribution. This process inherently allows for the propagation of uncertainty in a closed form, enabling the ensemble to capture and represent epistemic uncertainty arising from disagreements or variations across different models.
Bayes-optimal Coefficients
To determine the optimal coefficients $\bm{\pi}_{k}(\mathbf{x})$ for Gaussian convolution, we define an uncertainty-aware surrogate loss that reflects how closely the embeddings of positive pairs are expected to align. We then derive a convex optimization problem that identifies the coefficients minimizing this surrogate loss.
Text embeddings are often trained with contrastive losses, such as InfoNCE (Oord et al., 2018). For $\ell_{2}$-normalized features, the squared Euclidean distance relates directly to cosine similarity: $\|\mathbf{u}-\mathbf{v}\|^{2}=2(1-\mathbf{u}^{\top}\mathbf{v})$ (Wang & Isola, 2020). These losses also encourage alignment of positive pairs and uniformity on the unit hypersphere. Leveraging this connection, we adopt the squared loss as a principled surrogate for contrastive objectives.
Let $(\mathbf{x},\mathbf{x}^{\prime})$ be a positive pair. For each model $k$ we observe $\mathbf{z}_{k}(\mathbf{x})\sim\mathcal{N}(\bm{\mu}_{k}(\mathbf{x}),\bm{\Sigma}_{k}(\mathbf{x}))$ and $\mathbf{z}_{k}(\mathbf{x^{\prime}})\sim\mathcal{N}(\bm{\mu}_{k}(\mathbf{x^{\prime}}),\bm{\Sigma}_{k}(\mathbf{x^{\prime}}))$. We define the uncertainty-aware surrogate loss, a multi-task objective that aggregates model-specific embedding errors weighted by uncertainty-aware task coefficients, as:
$\mathcal{L}_{\mathrm{sur}}(\bm{\pi};\mathbf{x},\mathbf{x^{\prime}})$ $=\sum_{k=1}^{K}\bm{\pi}_{k}(\mathbf{x})\,\mathbb{E}_{\mathbf{z}_{k}(\mathbf{x})}\left[\|\mathbf{z}_{k}(\mathbf{x})-\mathbf{z}_{k}(\mathbf{x^{\prime}})\|^{2}\right]$
$=\sum_{k=1}^{K}\bm{\pi}_{k}(\mathbf{x})\left(\underbrace{\|\bm{\mu}_{k}(\mathbf{x})-\bm{\mu}_{k}(\mathbf{x^{\prime}})\|^{2}}_{\text{Fidelity}}+\underbrace{\operatorname{tr}(\bm{\Sigma}_{k}(\mathbf{x}))+\operatorname{tr}(\bm{\Sigma}_{k}(\mathbf{x^{\prime}}))}_{\text{Uncertainty}}\right).$ (2)
This surrogate loss can be decomposed into fidelity and epistemic uncertainty. While $\bm{\mu}_{k}(\mathbf{x})$ and $\bm{\Sigma}_{k}(\mathbf{x})$ are directly computable from the input $\mathbf{x}$, the corresponding terms $\bm{\mu}_{k}(\mathbf{x}^{\prime})$ and $\bm{\Sigma}_{k}(\mathbf{x}^{\prime})$ require evaluating the embedding model at the reference point $\mathbf{x}^{\prime}$. However, in our retrieval setting—aligned with common practice—document embeddings must be precomputed and stored in advance, independent of the incoming query $\mathbf{x}^{\prime}$. This practical constraint makes it infeasible to rely on any query-dependent terms during document embedding construction. As such, we approximate the surrogate loss by retaining only the components that depend on $\mathbf{x}$, leading to the following approximation:
$\min_{\bm{\pi}}\sum_{k=1}^{K}\bm{\pi}_{k}(\mathbf{x})\left(\operatorname{tr}(\bm{\Sigma}_{k}(\mathbf{x}))+\|\bm{\mu}_{k}(\mathbf{x})\|^{2}\right)\quad\text{s.t.}\;\sum_{k=1}^{K}\bm{\pi}_{k}(\mathbf{x})=\mathbf{1},\;\bm{\pi}_{k}(\mathbf{x})\geq 0.$ (3)
This is a convex optimization problem over the probability simplex, where each model is penalized based on both its uncertainty and the magnitude of its embedding. Its unique closed-form solution is the inverse-variance weighting:
$\bm{\pi}_{k}^{\star}(\mathbf{x})\approx\frac{1/\operatorname{tr}(\bm{\Sigma}_{k}(\mathbf{x}))}{\sum_{j=1}^{K}1/\operatorname{tr}(\bm{\Sigma}_{j}(\mathbf{x}))}.$ (4)
Eq. 4 generalized inverse-weighting scheme accounts for both epistemic uncertainty (via $\operatorname{tr}(\bm{\Sigma}_{k}(\mathbf{x}))$ and representation fidelity (via $\|\bm{\mu}_{k}(\mathbf{x})\|^{2}$), allowing the ensemble to favor embeddings that are simultaneously confident and well-aligned with the input. This optimal coefficient $\bm{\pi}^{*}_{k}(\mathbf{x})$ minimizes the uncertainty-aware surrogate loss in Eq. 3, maintaining consistency with the original objective. Full derivation appears in Appendix A.1.
3.3 Uncertainty-aware Similarity Estimation
Measuring distances between embeddings is fundamental to many downstream applications. While distributional distances like KL divergence (Kullback & Leibler, 1951) or Wasserstein distance (Chhachhi & Teng, 2023; Gelbrich, 1990) are theoretically well-founded, their computational cost is often prohibitive in large-scale settings. To address this, we propose a lightweight, principled similarity estimator.
Let the ensembled embeddings for $\mathbf{x}$ and $\mathbf{x^{\prime}}$ be $\mathbf{q}=\mathbf{z}(\mathbf{x})\sim\mathcal{N}(\bm{\mu}_{\mathbf{q}},\bm{\Sigma}_{\mathbf{q}})$ and $\mathbf{c}=\mathbf{z}(\mathbf{x^{\prime}})\sim\mathcal{N}(\bm{\mu}_{\mathbf{c}},\bm{\Sigma}_{\mathbf{c}})$, respectively. Our similarity measure $s$ is grounded in the properties of the base embedding models, which are typically trained with contrastive objectives on $\ell_{2}$-normalized outputs. This training process naturally results in mean vectors $\bm{\mu}_{\mathbf{q}}$, $\bm{\mu}_{\mathbf{c}}$ that are already near unit-norm, which in turn makes our small-variance assumption highly plausible. Along with it, we explicitly normalize the mean vectors, allowing us to approximate cosine similarity with the dot product, $s\approx\mathbf{q}^{\top}\mathbf{c}$. This operation preserves the distribution’s tractable Gaussian form, and the intrinsically small variance ensures the approximation is accurate. We then model the resulting dot product $s$ as a Gaussian via moment matching (Randone et al., 2024; Mallik & Sagias, 2011):
$s\sim\mathcal{N}(\mu_{s},\sigma_{s}^{2}),\quad\mu_{s}=\bm{\mu}_{\mathbf{q}}^{\top}\bm{\mu}_{\mathbf{c}},\quad\sigma_{s}^{2}=\bm{\mu}_{\mathbf{q}}^{\top}\bm{\Sigma}_{\mathbf{c}}\bm{\mu}_{\mathbf{q}}+\bm{\mu}_{\mathbf{c}}^{\top}\bm{\Sigma}_{\mathbf{q}}\bm{\mu}_{\mathbf{c}}+\mathrm{tr}(\bm{\Sigma}_{\mathbf{q}}\bm{\Sigma}_{\mathbf{c}}).$ (5)
To incorporate uncertainty, we adopt a probit approximation (Eschenhagen et al., 2021; Gibbs, 1998):
$\hat{s}\approx\frac{\mu_{s}}{\sqrt{1+\tfrac{\pi}{8}\sigma_{s}^{2}}}.$ (6)
The $\hat{s}$ incorporates predictive uncertainty in a lightweight manner without additional sampling. Beyond its computational efficiency, the estimator $\hat{s}$ is theoretically grounded, as it approximates the squared $2$-Wasserstein distance with provably bounded error, ensuring consistent ranking behavior.
Theorem 1 (Bounded approximation to the squared 2-Wasserstein distance).
Let $\mathbf{q}\sim\mathcal{N}(\bm{\mu}_{\mathbf{q}},\bm{\Sigma}_{\mathbf{q}})$ and $\mathbf{c}\sim\mathcal{N}(\bm{\mu}_{\mathbf{c}},\bm{\Sigma}_{\mathbf{c}})$ be two Gaussian embeddings with $\ell_{2}$-normalized mean vectors and diagonal covariances $\bm{\Sigma}_{\mathbf{q}}=\mathrm{diag}(\sigma_{\mathbf{q},1}^{2},\dots,\sigma_{\mathbf{q},d}^{2})$, $\bm{\Sigma}_{\mathbf{c}}=\mathrm{diag}(\sigma_{\mathbf{c},1}^{2},\dots,\sigma_{\mathbf{c},d}^{2})$. Assume the small-variance regime, where $\varepsilon:=\max_{i}\{\sigma_{\mathbf{q},i}^{2},\sigma_{\mathbf{c},i}^{2}\}<1$. Then, our similarity estimator $\hat{s}$ approximates the squared $2$-Wasserstein distance $W_{2}^{2}$ as:
$\hat{s}\;=\;1-\tfrac{1}{2}W_{2}^{2}+O(\varepsilon^{2}).$
Hence, ranking by $\hat{s}$ is equivalent to minimizing $W_{2}^{2}$, up to $O(\varepsilon^{2})$ error.
Our similarity estimator thus serves as a theoretically grounded surrogate for established distributional distances, achieving nearly identical ranking performance while being significantly more efficient. We provide a detailed derivation of the estimator in Appendix A.2, the proof for Theorem 1 in Appendix A.3, and empirical comparisons with alternative distances in Appendix B.2.
4 Experiments
Figure 3: Performance on MIRACL Subset across ensemble methods. The oracle represents the upper bound by selecting the best language-specific model per language. UEC achieves performances comparable to the oracle and even surpasses in some cases, with particularly strong gains in AUC@10.
4.1 MIRACL Subset
Setting
We first build a toy example with language-specialized models: E5 (Wang et al., 2022) (English/Russian), BGE-zh (Xiao et al., 2023) (Chinese), and MATRYOSHKA (Kusupati et al., 2024; Henderson et al., 2017) (Arabic). From the MIRACL Hard Negative dataset (Zhang et al., 2023), we sample 25 query-passage pairs each in English, Russian, Chinese, and Arabic. Baselines include uniform and weighted ensembles, along with an oracle upper bound that selects the best model per language. We report nDCG@10 and Recall@10 for retrieval and AUC@10 for uncertainty (Enevoldsen et al., 2025). Full experimental details are in Appendix B.3.
Figure 4: Heatmap of model-wise coefficients assigned by UEC per language. Each row corresponds to a language-specific input, and each column to an ensemble coefficient. UEC computes ensemble coefficients that are adaptively modulated by the uncertainty of each embedding.
Results
As shown in Figure 3, UEC consistently outperforms uniform and weighted ensembles. For retrieval metrics (nDCG@10, Recall@10), it achieves performance comparable to the oracle. UEC explicitly models embedding uncertainty and adaptively down-weights unreliable representations, leading to similarity scores that are better aligned with retrieval correctness. This enables UEC to surpass the oracle on the uncertainty metric AUC@10, highlighting that uncertainty-aware ensembling not only preserves strong retrieval performance but also delivers superior confidence calibration.
To qualitatively assess how UEC adapts ensemble weights via uncertainty, we visualize the averaged model-wise coefficients per language in Figure 4. The $4\times 3$ heatmap shows languages (rows: English, Russian, Chinese, Arabic) against ensembled models (columns). UEC assigns distinct patterns by language—for example, Arabic inputs emphasize Arabic embedding, while Chinese inputs favor Chinese embedding. It demonstrates that UEC effectively leverages embedding uncertainty to assign coefficients.
4.2 MMTEB
We further evaluate on a subset of MMTEB (Enevoldsen et al., 2025), focusing on three representative tasks: retrieval, classification, and Semantic Textual Similarity (STS). For each task, we carefully select datasets to cover a diverse range of domains and topics. Additional task results, which show performance consistent with the main results, are provided in the Appendix B.4.5. We use three SBERT-style base models—BGE (Xiao et al., 2023), E5 (Wang et al., 2022), and GTE (Li et al., 2023)—and a competitive multilingual baseline (GTE-MB (Zhang et al., 2024)). As a probabilistic embedding baseline, we adopt GroVE (Venkataramanan et al., 2025), applying it on top of the strongest individual model GTE-MB for fair comparison. To the best of our knowledge, this work presents the first systematic evaluation of model-merging (uniform/weighted, task arithmetic (Ilharco et al., 2022)) and ensemble (uniform/weighted) techniques for this benchmark. See Appendix B.4 for more experimental details.
4.2.1 Retrieval
Setting
We evaluate retrieval performance on five datasets: SCIDOCS (Cohan et al., 2020), LegalBenchCorporateLobbying (Guha et al., 2023), BelebeleRetrieval (Lovenia et al., 2024), WikipediaRetrievalMultilingual, and StackOverflowQA (Li et al., 2024). We use nDCG@10 and Recall@100 as retrieval metrics. In addition to standard retrieval metrics, we use AUC@10 (Enevoldsen et al., 2025) as a key quantitative measure to assess how well the model’s confidence scores are calibrated with its actual performance, directly evaluating the quality of its uncertainty estimation. We additionally compare Borda count (Emerson, 2013) as a representative rank aggregation technique, a natural strategy for retrieval tasks where each embedding model yields its own ranks.
Table 1: Comparison of retrieval performance. The best and second-best results per row are bolded and underlined. UEC achieves the highest average performance across all metrics, with particularly strong gains on the AUC@10 uncertainty metric, highlighting the efficiency for uncertainty calibration. (Borda=Borda count, Uni.=Uniform, Wt.=Weighted, TA=Task arithmetic).
Individual Models Borda Model Merging Ensemble Dataset / Metric BGE E5 GTE GTE-MB GroVE Uni. Wt. TA Uni. Wt. UEC SCIDOCS    nDCG@10 21.47 18.30 23.14 19.46 21.88 22.18 21.73 22.43 22.66 22.67 22.11 24.01    Recall@100 51.04 41.71 53.27 45.03 50.16 51.97 49.00 51.31 52.87 52.98 50.53 54.36    AUC@10 27.25 34.42 31.63 32.52 34.03 5.84 29.42 30.99 30.72 28.52 32.08 35.21 Lobbying    nDCG@10 90.42 91.54 91.81 90.40 90.91 91.23 92.35 92.32 90.91 91.79 92.15 93.56    Recall@100 99.70 100.00 99.70 99.70 99.70 99.70 100.00 99.70 99.41 100.0 100.0 100.0    AUC@10 55.56 58.24 55.02 54.62 57.81 32.83 57.49 53.17 53.18 52.07 55.79 60.04 Belebele    nDCG@10 93.64 94.91 93.76 94.09 94.14 96.26 94.60 94.86 93.87 95.25 93.64 95.82    Recall@100 99.44 99.55 99.77 99.66 99.44 99.44 99.66 99.66 99.44 99.66 99.55 99.77    AUC@10 83.83 84.78 86.18 82.53 86.57 69.17 86.39 85.11 86.03 82.99 83.34 88.72 Wikipedia    nDCG@10 92.17 93.32 92.71 91.55 92.71 93.12 91.88 93.08 92.99 93.32 93.84 94.24    Recall@100 99.80 99.93 99.93 99.80 99.86 99.55 99.86 99.86 99.80 99.86 99.86 99.93    AUC@10 59.84 61.06 61.57 62.08 62.36 49.26 58.61 60.60 62.62 62.41 62.01 64.25 StackOverflow    nDCG@10 79.93 87.14 80.54 91.17 87.76 85.36 78.93 81.31 80.55 85.76 88.71 90.26    Recall@100 96.59 97.54 97.64 98.79 97.54 98.09 96.28 97.44 96.54 98.09 98.09 99.40    AUC@10 69.59 84.00 68.28 88.61 85.06 17.14 69.69 73.89 71.27 75.51 82.43 89.83 Avg. nDCG@10 $\uparrow$ 75.52 77.04 76.39 77.33 77.48 77.11 75.90 75.69 76.19 77.76 78.49 79.58 Avg. Recall@100 $\uparrow$ 89.31 87.74 90.06 88.60 89.34 89.75 88.96 89.59 89.61 90.12 89.61 90.69 Avg. AUC@10 $\uparrow$ 54.65 64.70 60.74 64.07 65.16 34.85 60.52 60.75 60.76 60.30 63.13 67.61
Results
As shown in Table 1, UEC attains the highest average performance across both retrieval and uncertainty estimation, surpassing even the strong GTE-MB baseline by ensembling three weaker models. These results demonstrate that our uncertainty-aware ensemble framework effectively integrates heterogeneous embeddings, achieving robust and accurate retrieval across diverse datasets.
4.2.2 Classification
Setting
We evaluate on five classification datasets covering finance, law, poetry, multilingual intent, and social media: FinancialPhrasebankClassification (Malo et al., 2014), SwissJudgementClassification (Niklaus et al., 2021), PoemSentimentClassification (Sheng & Uthus, 2020), MassiveIntentClassification (FitzGerald et al., 2022), and TweetTopicSingleClassification (Antypas et al., 2022). Accuracy, F1 score, and AUROC are reported as evaluation metrics, where AUROC (Kuhn et al., 2023) serves as a key quantitative measure to assess how well the model estimates the degree of uncertainty in distinguishing between classes.
Table 2: Comparison of classification performance. The best and second-best results per row are bolded and underlined. UEC achieves the highest average scores, demonstrating robust performance across diverse classification tasks (Uni.=Uniform, Wt.=Weighted, TA=Task arithmetic).
Individual Models Model Merging Ensemble Dataset / Metric BGE E5 GTE GTE-MB GroVE Unif. Wt. TA Unif. Wt. UEC Financial    Accuracy 80.47 82.92 81.89 75.07 76.73 77.64 79.39 79.53 82.15 82.64 83.02    F1 78.38 79.90 79.10 73.50 75.65 75.53 77.85 77.93 79.76 80.18 80.30    AUROC 92.46 93.04 92.02 90.62 91.13 89.63 91.92 92.38 92.86 93.16 94.07 Judgement    Accuracy 57.16 58.58 56.37 55.45 55.44 56.82 55.90 55.44 57.09 57.71 58.36    F1 46.63 47.25 46.34 46.41 45.82 46.55 45.87 45.73 46.61 47.02 48.52    AUROC 50.48 50.22 50.65 50.67 50.48 50.13 50.24 50.37 50.43 50.51 51.02 Poem    Accuracy 51.92 53.07 47.88 56.92 53.13 50.28 49.90 51.82 52.30 52.01 56.81    F1 40.76 41.43 37.53 45.53 41.82 39.27 39.36 40.93 40.96 41.00 44.46    AUROC 56.70 57.38 53.41 62.21 57.31 55.13 54.36 55.76 56.71 56.27 61.39 Intent    Accuracy 73.46 68.05 65.09 75.19 74.29 72.95 74.07 75.33 68.78 68.31 77.08    F1 70.01 63.00 60.62 72.79 72.80 71.01 72.10 73.65 66.43 65.46 74.78    AUROC 66.74 64.54 63.09 67.23 66.98 64.07 67.15 67.02 65.26 65.11 67.91 Topic    Accuracy 71.75 69.96 71.97 70.99 71.24 71.18 70.59 70.53 73.60 73.47 74.20    F1 55.04 53.09 54.65 54.85 54.93 55.48 54.78 54.82 56.32 56.13 57.14    AUROC 87.02 87.83 88.54 89.76 87.95 88.64 88.12 88.33 90.07 90.12 90.73 Avg. Accuracy $\uparrow$ 66.95 66.52 64.64 66.73 66.16 65.77 65.97 66.53 66.79 66.83 68.89 Avg. F1 $\uparrow$ 58.16 56.93 55.65 58.62 58.20 57.57 57.99 58.61 58.02 57.96 61.04 Avg. AUROC $\uparrow$ 70.68 70.60 69.54 72.09 70.77 69.52 70.35 70.77 71.06 71.03 73.02
Results
As shown in Table 2, UEC attains the highest average accuracy, F1 score, and AUROC across five classification datasets. Notably, UEC also yields the best average AUROC, indicating that the model effectively captures the degree of uncertainty. These results demonstrate the effectiveness of UEC in classification scenarios where embedding-level uncertainty plays a crucial role.
4.2.3 Semantic Textual Similarity
Setting
We evaluate STS performance on 10 datasets, including STSBenchmark, FinParaSTS, SICK-R Marelli et al. (2014), STS22.v2 Chen et al. (2022), SemRel24STS (Ousidhoum et al., 2024), and STS12–STS17 (Agirre et al., 2012; 2013; Bandhakavi et al., 2014; Biçici, 2015; Cer et al., 2017) . Spearman correlation is used to measure alignment with human-annotated similarity scores.
Table 3: Comparison of STS performance. The best and second-best results per column are bolded and underlined. UEC achieves the highest average performance, ranking first on 8 out of the 10 datasets evaluated (Uni.=Uniform, Wt.=Weighted, TA=Task arithmetic).
Models STSB FinPara SICK-R SemRel24 STS12 STS13 STS14 STS15 STS17 STS22 Avg. Individual BGE 86.41 9.43 80.30 79.52 78.02 84.18 82.27 87.95 86.41 66.53 74.10 E5 85.47 15.99 78.39 81.37 73.49 82.99 80.44 88.18 88.89 66.51 74.17 GTE 85.73 12.95 78.85 77.79 75.70 85.72 81.51 88.80 87.88 67.65 74.26 GTE-MB 85.41 15.18 75.71 75.02 74.82 86.63 78.75 85.39 88.31 72.28 73.75 GroVE 85.93 11.53 80.02 78.99 76.26 83.33 79.03 85.27 88.89 66.98 73.62 Model Merging Uni. 86.78 11.75 79.85 79.49 75.43 86.79 82.07 89.12 88.71 68.42 74.84 Wt. 87.22 11.64 80.16 78.80 77.36 87.08 83.39 89.49 88.43 67.51 75.11 TA 86.83 10.22 80.25 78.61 78.40 85.85 83.18 88.67 87.41 66.31 74.57 Ensemble Uni. 87.44 12.03 80.73 80.24 78.00 85.90 83.62 89.07 88.21 68.15 75.34 Wt. 87.44 14.13 80.41 81.08 77.07 85.76 83.45 89.17 89.18 68.68 75.64 UEC 87.55 15.53 81.37 81.89 78.85 87.62 83.75 89.50 89.24 69.55 76.49
Results
As shown in Table 3, UEC achieves the highest average performance across the STS tasks. It ranks first on 8 out of the 10 datasets. These results demonstrate that our uncertainty-aware aggregation not only enhances robustness across diverse semantic similarity tasks but also outperforms existing deterministic and ensemble baselines.
4.3 Efficiency and Computational Characteristics
Table 4 compares the computational characteristics of UEC against common ensemble baselines. Three key capabilities are considered: automatic coefficient selection, data-wise coefficient estimation, and uncertainty-aware similarity estimation. Among all methods, only UEC supports all three features (), while others lack adaptability or require manual tuning (, -).
Table 4: Comparison of computational characteristics. UEC supports automatic coefficient selection and uncertainty-aware similarity while preserving efficiency. “Rel.” and “Comp.” denote relative time and complexity, respectively. Symbols: (Yes), (No), – (Not applicable).
Method Auto Coeff. Selection Data-wise Coeff. Uncertainty-aware Similarity Memory Complex. Time Rel. Complex. Uniform Ensemble - $\mathcal{O}(KD)$ 1.000 $\mathcal{O}(KD)$ Weighted Ensemble $\mathcal{O}(KD)$ 1.000 $\mathcal{O}(KD)$ Task Arithmetic $\mathcal{O}(KD)$ 1.000 $\mathcal{O}(KD)$ UEC (Ours) $\mathcal{O}(KD)$ 1.006 $\mathcal{O}(KD)$
Despite its added functionality, UEC introduces negligible overhead. It maintains the same asymptotic time and memory complexity $\mathcal{O}(KD)$, where $D$ is the embedding dimension. UEC increases only 0.6% in actual computation time for similarity estimation, demonstrating strong adaptivity without sacrificing efficiency and indicating its suitability for real-time deployment.
4.4 Ablation Study
Table 5: Ablation results for UEC’s components: uncertainty-aware similarity (Unc Sim) and adaptive convolution (Unc Conv). Removing either degrades performance, with the joint removal yielding the worst results.
nDCG@10 Recall@10 AUC@10 UEC 59.65% 80.07% 91.04% - Unc Sim 58.72% 78.13% 82.48% ($\downarrow$ 0.93%) ($\downarrow$1.94%) ($\downarrow$8.56%) - Unc Conv 48.45% 66.69% 10.30% ($\downarrow$ 11.20%) ($\downarrow$ 13.38%) ($\downarrow$ 80.74%) - Unc Sim, Conv 46.78% 62.66% 4.01% ($\downarrow$ 22.87%) ($\downarrow$ 17.41%) ($\downarrow$ 87.03%)
We perform an ablation study to evaluate each module’s contribution. Following the protocol in Section 4.1, we replace (i) the uncertainty-aware similarity with an uncalibrated one (- Unc Sim), (ii) the uncertainty-driven convolution coefficient with a uniform one (- Unc Conv), and (iii) both (- Unc Sim, Conv). Table 5 shows that removing any component noticeably drops performance, highlighting their individual and complementary contributions.
4.5 Challenging Case
Figure 5: Challenging case where only UEC correctly retrieves the positive passage, demonstrating its ability to leverage uncertainty.
Figure 5 shows a qualitative example from the setting in Section 4.1. Although the query is in English and E5 is expected to perform well, all individual models fail to retrieve the correct passage, which appears at ranks 12, 28, and 37. In contrast, UEC retrieves it at rank 6, thereby correctly retrieving the answer under both nDCG@10 and Recall@10 metrics, which are evaluated over the top 10 candidates. This example highlights UEC’s ability to handle model limitations by using uncertainty-aware ensembling and to generalize beyond language-specific encoders.
5 Conclusion and Discussion
We propose the Uncertainty-driven Embedding Convolution (UEC) framework, which converts pre-trained deterministic embeddings into probabilistic representations and adaptively ensembles them using estimated uncertainty, incorporating an uncertainty-aware similarity function. Experiments on diverse tasks show that UEC consistently outperforms strong baselines by leveraging complementary embedding strengths. While our study focuses on text embeddings and uses diagonal LA for uncertainty estimation, extending UEC to multimodal settings and exploring more scalable or advanced uncertainty methods remain promising future directions. Limitations include the need for embeddings of the same dimensionality and potential risks when applied to biased or incomplete data, underscoring the importance of responsible deployment with bias mitigation and transparent uncertainty communication. Overall, UEC enhances robustness and reliability in NLP systems, particularly in high-stakes domains.
Appendix A Derivation and Proof
A.1 Derivation of Bayes-optimal Weights for Uncertainty-driven Embedding Convolution
In this section, we provide a detailed derivation of the optimal ensemble coefficients used in the uncertainty-driven embedding convolution described in Section 3.2. Specifically, we aim to derive a closed-form solution for the optimal coefficients $\bm{\pi}_{k}(\mathbf{x})$ that minimize an uncertainty-aware surrogate loss reflecting the expected distance between positive pairs.
Problem Setting
Let there be $K$ independent embedding models. Each model $k\in\{1,\dots,K\}$ produces a probabilistic embedding $\mathbf{z}_{k}(\mathbf{x})\sim\mathcal{N}(\bm{\mu}_{k}(\mathbf{x}),\bm{\Sigma}_{k}(\mathbf{x}))$ for input $\mathbf{x}$, where $\bm{\mu}_{k}(\mathbf{x})\in\mathbb{R}^{d}$ is the mean and $\bm{\Sigma}_{k}(\mathbf{x})\in\mathbb{R}^{d\times d}$ is the covariance matrix. Similarly, we define $\mathbf{z}_{k}(\mathbf{x}^{\prime})\sim\mathcal{N}(\bm{\mu}_{k}(\mathbf{x}^{\prime}),\bm{\Sigma}_{k}(\mathbf{x}^{\prime}))$ for a positive example $\mathbf{x}^{\prime}$.
We define the ensembled embedding as a weighted combination of model-specific embeddings:
$\mathbf{z}(\mathbf{x})=\sum_{k=1}^{K}\bm{\pi}_{k}(\mathbf{x})\cdot\mathbf{z}_{k}(\mathbf{x}),\quad\text{with }\sum_{k=1}^{K}\bm{\pi}_{k}(\mathbf{x})=\mathbf{1},\quad\bm{\pi}_{k}(\mathbf{x})\geq 0.$
Surrogate Loss
Text embeddings are often trained with contrastive losses such as InfoNCE (Oord et al., 2018), which aim to bring semantically similar pairs closer while pushing dissimilar pairs apart. In the case of $\ell_{2}$-normalized features, the squared Euclidean distance is closely related to cosine similarity via the identity: $\|\mathbf{u}-\mathbf{v}\|^{2}=2(1-\mathbf{u}^{\top}\mathbf{v})$ (Wang & Isola, 2020). This connection reveals that minimizing squared distance effectively encourages high cosine similarity, aligning well with the objectives of contrastive learning.
Furthermore, recent theoretical analyses (Wang & Isola, 2020) show that contrastive objectives optimize two essential properties: (1) alignment, by pulling positive pairs together, and (2) uniformity, by encouraging embeddings to spread uniformly on the hypersphere. These principles are naturally supported by using the squared loss in the normalized feature space.
Leveraging this insight, we adopt the squared Euclidean loss as a principled surrogate for contrastive objectives. Unlike InfoNCE, which requires negative sampling and temperature scaling, the squared loss admits a simpler and deterministic formulation while still capturing the key inductive biases of alignment and uniformity. For a positive pair $(\mathbf{x},\mathbf{x}^{\prime})$, and for each embedding model $k$, we define the surrogate loss as the expected squared distance between probabilistic embeddings:
$\mathcal{L}_{\text{sur}}(\bm{\pi};\mathbf{x},\mathbf{x}^{\prime})=\sum_{k=1}^{K}\bm{\pi}_{k}(\mathbf{x})\cdot\mathbb{E}_{\mathbf{z}_{k}(\mathbf{x})}\left[\|\mathbf{z}_{k}(\mathbf{x})-\mathbf{z}_{k}(\mathbf{x}^{\prime})\|^{2}\right].$
Expected Distance Between Gaussians
To compute the surrogate loss, we evaluate the expected squared distance between two independent Gaussian embeddings:
$\mathbb{E}[\|\mathbf{z}_{k}(\mathbf{x})-\mathbf{z}_{k}(\mathbf{x}^{\prime})\|^{2}],$
where $\mathbf{z}_{k}(\mathbf{x})\sim\mathcal{N}(\bm{\mu}_{k}(\mathbf{x}),\bm{\Sigma}_{k}(\mathbf{x}))$ and $\mathbf{z}_{k}(\mathbf{x}^{\prime})\sim\mathcal{N}(\bm{\mu}_{k}(\mathbf{x}^{\prime}),\bm{\Sigma}_{k}(\mathbf{x}^{\prime}))$.
Using the identity $\|\mathbf{a}-\mathbf{b}\|^{2}=\|\mathbf{a}\|^{2}+\|\mathbf{b}\|^{2}-2\mathbf{a}^{\top}\mathbf{b}$, the expectation becomes:
$\mathbb{E}[\|\mathbf{z}_{k}(\mathbf{x})-\mathbf{z}_{k}(\mathbf{x}^{\prime})\|^{2}]$ $=\mathbb{E}[\|\mathbf{z}_{k}(\mathbf{x})\|^{2}]+\mathbb{E}[\|\mathbf{z}_{k}(\mathbf{x}^{\prime})\|^{2}]-2\mathbb{E}[\mathbf{z}_{k}(\mathbf{x})^{\top}\mathbf{z}_{k}(\mathbf{x}^{\prime})]$
$=\text{tr}(\bm{\Sigma}_{k}(\mathbf{x}))+\|\bm{\mu}_{k}(\mathbf{x})\|^{2}+\text{tr}(\bm{\Sigma}_{k}(\mathbf{x}^{\prime}))+\|\bm{\mu}_{k}(\mathbf{x}^{\prime})\|^{2}-2\bm{\mu}_{k}(\mathbf{x})^{\top}\bm{\mu}_{k}(\mathbf{x}^{\prime}),$
where we used two standard results for Gaussian vectors:
  • $\mathbb{E}[\|\mathbf{z}\|^{2}]=\text{tr}(\bm{\Sigma})+\|\bm{\mu}\|^{2}$, and
  • $\mathbb{E}[\mathbf{z}^{\top}\mathbf{z}^{\prime}]=\mu^{\top}\mu^{\prime}$ for independent $\mathbf{z}\sim\mathcal{N}(\bm{\mu},\bm{\Sigma}),\ \mathbf{z}^{\prime}\sim\mathcal{N}(\bm{\mu}^{\prime},\bm{\Sigma}^{\prime})$.
Combining these yields:
$\mathbb{E}[\|\mathbf{z}_{k}(\mathbf{x})-\mathbf{z}_{k}(\mathbf{x}^{\prime})\|^{2}]=\|\bm{\mu}_{k}(\mathbf{x})-\bm{\mu}_{k}(\mathbf{x}^{\prime})\|^{2}+\text{tr}(\bm{\Sigma}_{k}(\mathbf{x}))+\text{tr}(\bm{\Sigma}_{k}(\mathbf{x}^{\prime})).$
This decomposition reveals two key components:
  • The fidelity term $\|\bm{\mu}_{k}(\mathbf{x})-\bm{\mu}_{k}(\mathbf{x}^{\prime})\|^{2}$, which measures semantic alignment between the means of the embeddings.
  • The uncertainty term $\text{tr}(\bm{\Sigma}_{k}(\mathbf{x}))+\text{tr}(\bm{\Sigma}_{k}(\mathbf{x}^{\prime}))$, which reflects how uncertain the model is about each input.
This result plays a central role in our surrogate loss formulation, as it enables UEC to quantify both how close embeddings are (fidelity) and how reliable they are (uncertainty), using closed-form computations under the Gaussian assumption.
Tractable Approximation
Directly evaluating the full surrogate loss requires computing both the fidelity term $\|\bm{\mu}_{k}(\mathbf{x})-\bm{\mu}_{k}(\mathbf{x}^{\prime})\|^{2}$ and the uncertainty terms $\operatorname{tr}(\bm{\Sigma}_{k}(\mathbf{x}))$ and $\operatorname{tr}(\bm{\Sigma}_{k}(\mathbf{x}^{\prime}))$ for every positive pair $(\mathbf{x},\mathbf{x}^{\prime})$. However, in many practical settings, this becomes computationally prohibitive, as document embeddings (corresponding to $\mathbf{x}^{\prime}$) are typically precomputed and stored in advance. This design choice is essential for scalable retrieval systems, which need to index millions of documents offline and later perform fast matching against online queries $\mathbf{x}$ at inference time.
Consequently, the embedding $\bm{\mu}_{k}(\mathbf{x}^{\prime})$ and its associated uncertainty $\bm{\Sigma}_{k}(\mathbf{x}^{\prime})$ are fixed and query-independent at runtime. Moreover, the query $\mathbf{x}$ is not known during document encoding, so it is infeasible to compute any query-dependent quantities during offline document embedding. To ensure consistency with this retrieval pipeline, we simplify the surrogate loss by retaining only the components that are computable solely from the input $\mathbf{x}$ (i.e., the query):
$\min_{\pi}\sum_{k=1}^{K}\bm{\pi}_{k}(\mathbf{x})\left(\operatorname{tr}(\bm{\Sigma}_{k}(\mathbf{x}))+\|\bm{\mu}_{k}(\mathbf{x})\|^{2}\right)\quad\text{s.t.}\;\sum_{k=1}^{K}\bm{\pi}_{k}(\mathbf{x})=\mathbf{1},\;\bm{\pi}_{k}(\mathbf{x})\geq 0.$ (7)
This formulation preserves the uncertainty-awareness of our surrogate loss while adhering to the architectural constraint of precomputed document embeddings. It also allows for efficient, query-time optimization of the ensemble coefficients $\bm{\pi}_{k}(\mathbf{x})$ based only on the statistics of the query embedding.
Closed-form Solution
The optimization problem defined above (Eq. 7) is convex over the probability simplex and has a linear objective. The objective is minimized when more weight is assigned to components with lower values of the cost term $\mathbf{c}_{k}:=\operatorname{tr}(\bm{\Sigma}_{k}(\mathbf{x}))+\|\bm{\mu}_{k}(\mathbf{x})\|^{2}$.
We derive the solution using the method of Lagrange multipliers. The Lagrangian is:
$\mathcal{L}(\bm{\pi},\bm{\lambda})=\sum_{k=1}^{K}\bm{\pi}_{k}\mathbf{c}_{k}-\bm{\lambda}\left(\sum_{k=1}^{K}\bm{\pi}_{k}-1\right),$
where $\bm{\lambda}$ is the Lagrange multiplier. Taking the derivative with respect to $\bm{\pi}_{k}$ and setting to zero gives:
$\frac{\partial\mathcal{L}}{\partial\bm{\pi}_{k}}=\mathbf{c}_{k}-\bm{\lambda}=0\quad\Rightarrow\quad\bm{\lambda}=\mathbf{c}_{k}\quad\forall k\ \text{ s.t. }\ \bm{\pi}_{k}>0.$
This condition implies that the minimum is achieved when $\bm{\pi}_{k}$ is inversely proportional to $\mathbf{c}_{k}$. More formally, the optimal solution is:
$\bm{\pi}_{k}^{\star}(\mathbf{x})=\frac{\mathbf{c}_{k}^{-1}}{\sum_{j=1}^{K}\mathbf{c}_{j}^{-1}}=\frac{\left(\operatorname{tr}(\bm{\Sigma}_{k}(\mathbf{x}))+\|\bm{\mu}_{k}(\mathbf{x})\|^{2}\right)^{-1}}{\sum_{j=1}^{K}\left(\operatorname{tr}(\bm{\Sigma}_{j}(\mathbf{x}))+\|\bm{\mu}_{j}(\mathbf{x})\|^{2}\right)^{-1}}.$ (8)
In practice, the embeddings $\bm{\mu}_{k}(\mathbf{x})$ are typically $\ell_{2}$-normalized, which implies that $\|\bm{\mu}_{k}(\mathbf{x})\|^{2}\approx 1$ for all $k$. Therefore, we approximate Eq 8 as:
$\bm{\pi}_{k}^{\star}(\mathbf{x})\approx\frac{1/\operatorname{tr}(\bm{\Sigma}_{k}(\mathbf{x}))}{\sum_{j=1}^{K}1/\operatorname{tr}(\bm{\Sigma}_{j}(\mathbf{x}))},$
by treating the constant $\|\bm{\mu}_{k}(\mathbf{x})\|^{2}$ term as uniform across all models. This simplification preserves the desirable behavior of assigning higher weights to models with lower predictive uncertainty, while significantly reducing computational overhead.
Interpretation
This optimal solution is Bayes-optimal under the surrogate loss, as it minimizes the expected discrepancy between the ensembled embedding and its positive reference under Gaussian uncertainty. The form of the solution resembles that of precision-weighted model averaging, often seen in Bayesian linear models, where less uncertain models contribute more to the final prediction.
A.2 Derivation of Uncertainty-aware Similarity Estimation
In this section, we provide a detailed derivation of the uncertainty-aware similarity estimation method introduced in Section 3.3. When embeddings are modeled as multivariate Gaussian distributions, the similarity between them becomes a random variable. We aim to derive the distribution of this similarity score and propose a calibrated approximation that accounts for predictive uncertainty.
Problem Setup
Let $\mathbf{q}=\mathbf{z}(\mathbf{x})\sim\mathcal{N}(\bm{\mu}_{\mathbf{q}},\bm{\Sigma}_{\mathbf{q}})$ and $\mathbf{c}=\mathbf{z}(\mathbf{x}^{\prime})\sim\mathcal{N}(\bm{\mu}_{\mathbf{c}},\bm{\Sigma}_{\mathbf{c}})$ denote two ensembled probabilistic embeddings generated by the proposed UEC method. Each embedding is represented as a Gaussian distribution with a mean vector $\bm{\mu}\in\mathbb{R}^{d}$ and a covariance matrix $\bm{\Sigma}\in\mathbb{R}^{d\times d}$.
We define the cosine similarity between $\mathbf{q}$ and $\mathbf{c}$ as:
$s:=\cos(\mathbf{q},\mathbf{c})=\frac{\mathbf{q}^{\top}\mathbf{c}}{\|\mathbf{q}\|\cdot\|\mathbf{c}\|},$
which itself is a random variable, since both $\mathbf{q}$ and $\mathbf{c}$ are random. In our implementation, both $\mathbf{q}$ and $\mathbf{c}$ are $\ell_{2}$-normalized in expectation (i.e., $\|\bm{\mu}_{\mathbf{q}}\|\approx\|\bm{\mu}_{\mathbf{c}}\|\approx 1$), and their variances are relatively small. Therefore, we simplify cosine similarity to the inner product:
$s\approx\mathbf{q}^{\top}\mathbf{c},$
as is common in prior work (Mallik & Sagias, 2011; Randone et al., 2024).
Distribution of the Inner Product via Moment Matching
To characterize the distribution of $s=\mathbf{q}^{\top}\mathbf{c}$, we approximate it as a Gaussian random variable using moment-matching (Randone et al., 2024). Since $\mathbf{q}$ and $\mathbf{c}$ are independent Gaussian vectors, their inner product $s$ follows an approximately Gaussian distribution whose mean and variance can be computed in closed form.
Mean By linearity of expectation:
$\mu_{s}:=\mathbb{E}[\mathbf{q}^{\top}\mathbf{c}]=\mathbb{E}[\mathbf{q}]^{\top}\mathbb{E}[\mathbf{c}]=\bm{\mu}_{\mathbf{q}}^{\top}\bm{\mu}_{\mathbf{c}}.$ (9)
Variance Using the law of total variance:
$\mathrm{Var}(\mathbf{q}^{\top}\mathbf{c})$ $=\mathbb{E}_{\mathbf{q}}\left[\mathrm{Var}(\mathbf{q}^{\top}\mathbf{c}\mid\mathbf{q})\right]+\mathrm{Var}_{\mathbf{q}}\left(\mathbb{E}[\mathbf{q}^{\top}\mathbf{c}\mid\mathbf{q}]\right)$
$=\mathbb{E}_{\mathbf{q}}[\mathbf{q}^{\top}\bm{\Sigma}_{\mathbf{c}}\mathbf{q}]+\mathrm{Var}_{\mathbf{q}}(\mathbf{q}^{\top}\bm{\mu}_{\mathbf{c}}),$
where assume $\mathbf{q}$ and $\mathbf{c}$ are independent.
We evaluate each term as follows. For the first term, using the identity $\mathbb{E}[\mathbf{x}^{\top}\mathbf{A}\mathbf{x}]=\mathrm{tr}(\mathbf{A}\bm{\Sigma})+\bm{\mu}^{\top}\mathbf{A}\bm{\mu}$ for $\mathbf{x}\sim\mathcal{N}(\bm{\mu},\bm{\Sigma})$ (Shao, 2008):
$\mathbb{E}_{\mathbf{q}}[\mathbf{q}^{\top}\bm{\Sigma}_{\mathbf{c}}\mathbf{q}]=\mathrm{tr}(\bm{\Sigma}_{\mathbf{c}}\bm{\Sigma}_{\mathbf{q}})+\bm{\mu}_{\mathbf{q}}^{\top}\bm{\Sigma}_{\mathbf{c}}\bm{\mu}_{\mathbf{q}}.$
For the second term, using the variance of a linear form in a Gaussian:
$\mathrm{Var}_{\mathbf{q}}(\mathbf{q}^{\top}\bm{\mu}_{\mathbf{c}})=\bm{\mu}_{\mathbf{c}}^{\top}\bm{\Sigma}_{\mathbf{q}}\bm{\mu}_{\mathbf{c}}.$
Combining both, we obtain:
$\sigma_{s}^{2}:=\mathrm{Var}(\mathbf{q}^{\top}\mathbf{c})=\bm{\mu}_{\mathbf{q}}^{\top}\bm{\Sigma}_{\mathbf{c}}\bm{\mu}_{\mathbf{q}}+\bm{\mu}_{\mathbf{c}}^{\top}\bm{\Sigma}_{\mathbf{q}}\bm{\mu}_{\mathbf{c}}+\mathrm{tr}(\bm{\Sigma}_{\mathbf{q}}\bm{\Sigma}_{\mathbf{c}}).$
Therefore, under moment-matching, we approximate the distribution of similarity as:
$s=\mathbf{q}^{\top}\mathbf{c}\sim\mathcal{N}(\mu_{s},\sigma_{s}^{2}).$
Probit-based Calibration of Similarity Scores
The similarity distribution provides a way to quantify uncertainty, but we still need a scalar value to rank candidates. To account for this uncertainty in a principled yet computationally efficient way, we apply a probit-based calibration to the mean similarity score:
$\hat{s}\approx\frac{\mu_{s}}{\sqrt{1+\frac{\pi}{8}\sigma_{s}^{2}}}.$ (10)
This formula arises from approximating the expectation of a sigmoid over a Gaussian random variable (Gibbs, 1998; Eschenhagen et al., 2021), i.e., $\mathbb{E}[\sigma(\mathbf{z})]\approx\sigma\left(\mu/\sqrt{1+\frac{\pi}{8}\sigma^{2}}\right)$ when $z\sim\mathcal{N}(\mu,\sigma^{2})$ and $\sigma(\cdot)$ is the logistic sigmoid function.
Interpretation and Robustness
The numerator $\mu_{s}$ captures alignment between the means, while the denominator serves as a soft penalty for high uncertainty. As $\sigma_{s}^{2}$ increases, the similarity score $\hat{s}$ is downscaled, reflecting reduced confidence in the similarity estimate. This makes the scoring robust in the presence of model disagreement, variance, or noise, particularly useful in retrieval settings.
This uncertainty-aware formulation preserves the computational simplicity of cosine similarity while incorporating predictive variance in a principled manner, enabling better performance under high uncertainty.
A.3 Derivation of Theorem 1
We prove that the proposed estimator $\hat{s}$ provides a bounded approximation to the squared $2$-Wasserstein distance $W_{2}^{2}$ under diagonal-covariance assumptions, and we extend this bound to the Jeffreys divergence $J$ under additional regularity assumptions.
Recall that
$\mathbf{q}=\mathbf{z}(\mathbf{x})\sim\mathcal{N}(\bm{\mu}_{\mathbf{q}},\bm{\Sigma}_{\mathbf{q}}),\qquad\mathbf{c}=\mathbf{z}(\mathbf{x}^{\prime})\sim\mathcal{N}(\bm{\mu}_{\mathbf{c}},\bm{\Sigma}_{\mathbf{c}}),$
with $\ell_{2}$-normalized means $\|\bm{\mu}_{\mathbf{q}}\|=\|\bm{\mu}_{\mathbf{c}}\|=1$, and diagonal covariances
$\bm{\Sigma}_{\mathbf{q}}=\mathrm{diag}(\sigma_{\mathbf{q},1}^{2},\dots,\sigma_{\mathbf{q},d}^{2}),\qquad\bm{\Sigma}_{\mathbf{c}}=\mathrm{diag}(\sigma_{\mathbf{c},1}^{2},\dots,\sigma_{\mathbf{c},d}^{2}).$
We define the per-coordinate maximum variance
$\varepsilon:=\max_{i}\{\sigma_{\mathbf{q},i}^{2},\sigma_{\mathbf{c},i}^{2}\}.$
A.3.1 Preliminary
In Section 3.3 the dot product $s=\mathbf{q}^{\top}\mathbf{c}$ was approximated by a Gaussian with mean and variance
$\mu_{s}=\bm{\mu}_{\mathbf{q}}^{\top}\bm{\mu}_{\mathbf{c}},\qquad\sigma_{s}^{2}=\bm{\mu}_{\mathbf{q}}^{\top}\bm{\Sigma}_{\mathbf{c}}\bm{\mu}_{\mathbf{q}}+\bm{\mu}_{\mathbf{c}}^{\top}\bm{\Sigma}_{\mathbf{q}}\bm{\mu}_{\mathbf{c}}+\operatorname{tr}(\bm{\Sigma}_{\mathbf{q}}\bm{\Sigma}_{\mathbf{c}}).$
The probit-calibrated similarity estimator is then approximated as
$\hat{s}\;\approx\;\frac{\bm{\mu}_{\mathbf{q}}^{\top}\bm{\mu}_{\mathbf{c}}}{\sqrt{1+\tfrac{\pi}{8}\sigma_{s}^{2}}}.$
Using $\|\bm{\mu}_{\mathbf{q}}\|=\|\bm{\mu}_{\mathbf{c}}\|=1$, the mean term admits the exact identity
[…]
A.3.2 Approximation of Uncertainty-aware Similarity Estimation
Applying the Taylor expansion $(1+x)^{-1/2}=1-\tfrac{1}{2}x+O(x^{2})$ for small $x$, we linearize the denominator:
$\frac{1}{\sqrt{1+\tfrac{\pi}{8}\sigma_{s}^{2}}}\;=\;1-\tfrac{\pi}{16}\sigma_{s}^{2}+O(\sigma_{s}^{4}).$ (12)
Substituting Eq. 11 and Eq. 12 yields
$\hat{s}$ $\approx\Big(1-\tfrac{1}{2}\|\bm{\mu}_{\mathbf{q}}-\bm{\mu}_{\mathbf{c}}\|^{2}\Big)\Big(1-\tfrac{\pi}{16}\sigma_{s}^{2}\Big)+O(\sigma_{s}^{4})$
$=1-\tfrac{1}{2}\|\bm{\mu}_{\mathbf{q}}-\bm{\mu}_{\mathbf{c}}\|^{2}-\tfrac{\pi}{16}\sigma_{s}^{2}+O(\sigma_{s}^{4}).$
Rearranging and discarding the additive constant 1 which does not affect ranking gives the expansion
$\hat{s}\;=\;\tfrac{1}{2}\|\bm{\mu}_{\mathbf{q}}-\bm{\mu}_{\mathbf{c}}\|^{2}\;+\;\tfrac{\pi}{16}\sigma_{s}^{2}\;+\;O(\sigma_{s}^{4}).$ (13)
A.3.3 Comparison with squared 2-Wasserstein distance
We now compare the expansion Eq. 13 with the closed form of the squared 2-Wasserstein distance for diagonal covariances:
$W_{2}^{2}\;=\;\|\bm{\mu}_{\mathbf{q}}-\bm{\mu}_{\mathbf{c}}\|^{2}\;+\;\sum_{i=1}^{d}(\sigma_{\mathbf{q},i}-\sigma_{\mathbf{c},i})^{2}.$ (14)
It is useful to align affine factors so that the leading mean term cancels: using Eq. 13 we have
$2(1-\hat{s})\;=\;\|\bm{\mu}_{\mathbf{q}}-\bm{\mu}_{\mathbf{c}}\|^{2}+\tfrac{\pi}{8}\sigma_{s}^{2}+2O(\sigma_{s}^{4}).$ (15)
Define the residual (difference) between $W_{2}^{2}$ and the affine-aligned estimator:
[…]
Substituting Eq. 14 and the aligned form gives
$\Delta$ $=\bigg(\|\bm{\mu}_{\mathbf{q}}-\bm{\mu}_{\mathbf{c}}\|^{2}+\sum_{i=1}^{d}(\sigma_{\mathbf{q},i}-\sigma_{\mathbf{c},i})^{2}\bigg)-\bigg(\|\bm{\mu}_{\mathbf{q}}-\bm{\mu}_{\mathbf{c}}\|^{2}+\tfrac{\pi}{8}\sigma_{s}^{2}+2O(\sigma_{s}^{4})\bigg)$
$=\sum_{i=1}^{d}(\sigma_{\mathbf{q},i}-\sigma_{\mathbf{c},i})^{2}-\tfrac{\pi}{8}\sigma_{s}^{2}-2O(\sigma_{s}^{4}).$ (17)
Thus the mean separation term is exactly cancelled by aligning 2(1s^)2(1-$\hat${s}) with $W_{2}^{2}$; the residual $\Delta$ depends only on variance-related terms and the higher-order remainder $O(\sigma_{s}^{4})$. The task is to bound $\Delta$.
Elementwise Expansion
Write each term elementwise:
$\sum_{i=1}^{d}(\sigma_{\mathbf{q},i}-\sigma_{\mathbf{c},i})^{2}$ $=\sum_{i=1}^{d}(\sigma_{\mathbf{q},i}^{2}+\sigma_{\mathbf{c},i}^{2}-2\sigma_{\mathbf{q},i}\sigma_{\mathbf{c},i}),$
$\sigma_{s}^{2}$ $=\sum_{i=1}^{d}\big(\sigma_{\mathbf{c},i}^{2}\mu_{\mathbf{q},i}^{2}+\sigma_{\mathbf{q},i}^{2}\mu_{\mathbf{c},i}^{2}+\sigma_{\mathbf{q},i}^{2}\sigma_{\mathbf{c},i}^{2}\big).$
Substituting into Eq. 17 yields
$\Delta=\sum_{i=1}^{d}\Big[(\sigma_{\mathbf{q},i}^{2}+\sigma_{\mathbf{c},i}^{2}-2\sigma_{\mathbf{q},i}\sigma_{\mathbf{c},i})-\tfrac{\pi}{8}\big(\sigma_{\mathbf{}c,i}^{2}\mu_{\mathbf{q},i}^{2}+\sigma_{\mathbf{q},i}^{2}\mu_{\mathbf{c},i}^{2}+\sigma_{\mathbf{q},i}^{2}\sigma_{\mathbf{c},i}^{2}\big)\Big]-2O(\sigma_{s}^{4}).$ (18)
General Bound Without Extra Scaling Assumptions
Using only the trivial bounds $0\leq\mu_{(\cdot),i}^{2}\leq 1$ and $\sigma_{(\cdot),i}^{2}\leq\varepsilon$, we obtain the estimate
$\big|\Delta\big|\leq\sum_{i=1}^{d}\big(\sigma_{\mathbf{q},i}^{2}+\sigma_{\mathbf{c},i}^{2}+\tfrac{\pi}{8}(\sigma_{\mathbf{c},i}^{2}+\sigma_{\mathbf{q},i}^{2}+\sigma_{\mathbf{q},i}^{2}\sigma_{\mathbf{c},i}^{2})\big)+2|O(\sigma_{s}^{4})|$
$\qquad\leq C_{0}\big(\operatorname{tr}(\bm{\Sigma}_{\mathbf{q}})+\operatorname{tr}(\bm{\Sigma}_{\mathbf{c}})+\operatorname{tr}(\bm{\Sigma}_{\mathbf{q}}\bm{\Sigma}_{\mathbf{c}})\big)+O(\sigma_{s}^{4}),$
for some numerical constant $C_{0}$ (e.g. depending on $\bm{\pi}$). Since $\operatorname{tr}(\bm{\Sigma}_{\mathbf{q}}\bm{\Sigma}_{\mathbf{c}})=O(\varepsilon^{2})$ and $\operatorname{tr}(\bm{\Sigma}_{(\cdot)})=O(\varepsilon)$, the coarse bound yields $\Delta=O(\varepsilon)$ in general. This already justifies the statement that 2(1s^)2(1-$\hat${s}) approximates $W_{2}^{2}$ up to variance-controlled error; however the bound is not yet the tight $O(\varepsilon^{2})$ form.
Tighter Bound Under Small-variance Scaling
To obtain the strong $O(\varepsilon^{2})$ residual we introduce additional mild scaling assumptions which are natural in the embedding-ensemble / small-uncertainty regime:
Assumption 1 (Small-variance trace scaling).
$\operatorname{tr}(\bm{\Sigma}_{\mathbf{q}})=O(\varepsilon),\qquad\operatorname{tr}(\bm{\Sigma}_{\mathbf{c}})=O(\varepsilon).$
Assumption 2 (Per-dimension dispersion).
The per-coordinate variances are dispersed across dimensions, i.e.
$\max_{i}\{\sigma_{\mathbf{q},i}^{2},\sigma_{\mathbf{c},i}^{2}\}\leq\frac{\varepsilon}{d}.$
Both assumptions are mild in practice when the posterior variances of embedding coordinates are small and not concentrated on a few coordinates (a common situation for high-dimensional embeddings).
Elementwise Cancellation Argument
Under Assumptions 12 we refine the termwise bounds in Eq. 18. Since $\mu_{\mathbf{q},i}^{2}\leq 1$ and Assumption 2 yields $\sigma_{(\cdot),i}^{2}\leq\varepsilon/d$, each element in the bracket of Eq. 18 is
$(\sigma_{\mathbf{q},i}^{2}+\sigma_{\mathbf{c},i}^{2}-2\sigma_{\mathbf{q},i}\sigma_{\mathbf{c},i})-\tfrac{\pi}{8}(\sigma_{\mathbf{c},i}^{2}\mu_{\mathbf{q},i}^{2}+\sigma_{\mathbf{q},i}^{2}\mu_{\mathbf{c},i}^{2}+\sigma_{\mathbf{q},i}^{2}\sigma_{\mathbf{c},i}^{2})=O\!\left(\frac{\varepsilon^{2}}{d^{2}}\right).$
Summing over $i=1,\dots,d$ yields
$\sum_{i=1}^{d}O\!\left(\frac{\varepsilon^{2}}{d^{2}}\right)=O(\varepsilon^{2}).$
Moreover, the remainder term $O(\sigma_{s}^{4})$ is $O(\varepsilon^{2})$ under assumptions. Hence $\Delta=O(\varepsilon^{2})$. This proves the quantitative bound claimed in Theorem 1:
$\hat{s}\;=\;1-\tfrac{1}{2}W_{2}^{2}+O(\varepsilon^{2}).$
A.3.4 Comparison with Jeffreys divergence
We now show how a similar first-order argument extends to the Jeffreys divergence
$J\big(\mathcal{N}(\bm{\mu}_{\mathbf{q}},\bm{\Sigma}_{\mathbf{q}}),\mathcal{N}(\bm{\mu}_{\mathbf{c}},\bm{\Sigma}_{\mathbf{c}})\big):=D_{\mathrm{KL}}(\mathcal{N}(\bm{\mu}_{\mathbf{q}},\bm{\Sigma}_{\mathbf{q}})\|\mathcal{N}(\bm{\mu}_{\mathbf{c}},\bm{\Sigma}_{\mathbf{c}}))+D_{\mathrm{KL}}(\mathcal{N}(\bm{\mu}_{\mathbf{c}},\bm{\Sigma}_{\mathbf{c}})\|\mathcal{N}(\bm{\mu}_{\mathbf{q}},\bm{\Sigma}_{\mathbf{q}})),$
which for diagonal covariances admits the closed form
$J=\tfrac{1}{2}(\bm{\mu}_{\mathbf{c}}-\bm{\mu}_{\mathbf{q}})^{\top}(\bm{\Sigma}_{\mathbf{c}}^{-1}+\bm{\Sigma}_{\mathbf{q}}^{-1})(\bm{\mu}_{\mathbf{c}}-\bm{\mu}_{\mathbf{q}})\;+\;\tfrac{1}{2}\sum_{i=1}^{d}\Big(\tfrac{\sigma_{\mathbf{q},i}^{2}}{\sigma_{\mathbf{c},i}^{2}}+\tfrac{\sigma_{\mathbf{c},i}^{2}}{\sigma_{\mathbf{q},i}^{2}}-2\Big).$ (19)
Unlike the Wasserstein case, the Jeffreys mean-term involves inverse covariances. To linearize these terms we need a regularity assumption that the diagonal variances of $\mathbf{q}$ and $\mathbf{c}$ are not only small but also close to a common positive baseline. Specifically, we make the following:
Assumption 3 (Common-scale regularity).
There exists a baseline $\tau>0$ and constants $K,\kappa^{\prime}>0$ such that for all $i$,
$|\sigma_{\mathbf{q},i}^{2}-\tau|\leq K\varepsilon,\qquad|\sigma_{\mathbf{c},i}^{2}-\tau|\leq K\varepsilon,$
and $\tau$ is bounded away from zero (i.e. $\tau\geq\tau_{\min}>0$). Moreover, $\operatorname{tr}(\bm{\Sigma}_{\mathbf{q}}),\operatorname{tr}(\bm{\Sigma}_{\mathbf{c}})=O(\varepsilon)$ as in Assumption 1.
Assumption 3 asserts that each diagonal entry is a perturbation of a common (nonzero) scale $\tau$; the smallness parameter is the same $\varepsilon$ used above. Under this assumption we can Taylor-expand reciprocal diagonal entries.
Expansion of Inverse Covariances
For each $i$, write
$\sigma_{\mathbf{q},i}^{2}=\tau+\delta_{\mathbf{q},i},\qquad\sigma_{\mathbf{c},i}^{2}=\tau+\delta_{\mathbf{c},i},\qquad\text{with }|\delta_{(\cdot),i}|\leq K\varepsilon.$
Then, using the scalar expansion $(\tau+\delta)^{-1}=\tau^{-1}-\tau^{-2}\delta+O(\delta^{2})$, we obtain
$\bm{\Sigma}_{\mathbf{q}}^{-1}$ $=\tau^{-1}\mathbf{I}-\tau^{-2}\mathrm{diag}(\delta_{q,1},\dots,\delta_{q,d})+O(\varepsilon^{2}),$
$\bm{\Sigma}_{\mathbf{c}}^{-1}$ $=\tau^{-1}\mathbf{I}-\tau^{-2}\mathrm{diag}(\delta_{c,1},\dots,\delta_{c,d})+O(\varepsilon^{2}).$ (20)
Hence
$\bm{\Sigma}_{\mathbf{c}}^{-1}+\bm{\Sigma}_{\mathbf{q}}^{-1}=2\tau^{-1}\mathbf{I}-\tau^{-2}\mathrm{diag}(\delta_{q,1}+\delta_{c,1},\dots,\delta_{q,d}+\delta_{c,d})+O(\varepsilon^{2}).$
Mean Term in $J$
Substitute Eq. 20 into the mean quadratic term of Eq. 19:
$\tfrac{1}{2}(\bm{\mu}_{\mathbf{c}}-\bm{\mu}_{\mathbf{q}})^{\top}(\bm{\Sigma}_{\mathbf{c}}^{-1}+\bm{\Sigma}_{\mathbf{q}}^{-1})(\bm{\mu}_{\mathbf{c}}-\bm{\mu}_{\mathbf{q}})$ $=\tfrac{1}{2}(\bm{\mu}_{\mathbf{c}}-\bm{\mu}_{\mathbf{q}})^{\top}\big(2\tau^{-1}\mathbf{I}\big)(\bm{\mu}_{\mathbf{c}}-\bm{\mu}_{\mathbf{q}})$
$\qquad-\tfrac{1}{2}(\bm{\mu}_{\mathbf{c}}-\bm{\mu}_{\mathbf{q}})^{\top}\tau^{-2}\mathrm{diag}(\delta_{\mathbf{q},i}+\delta_{\mathbf{c},i})(\bm{\mu}_{\mathbf{c}}-\bm{\mu}_{\mathbf{q}})+O(\varepsilon^{2})$
$=\tau^{-1}\cdot\tfrac{1}{2}\|\bm{\mu}_{\mathbf{q}}-\bm{\mu}_{\mathbf{c}}\|^{2}+O(\varepsilon/d)+O(\varepsilon^{2}).$ (21)
Variance-only Term in $J$
Consider the variance-only summation in Eq. 19:
$V_{J}:=\tfrac{1}{2}\sum_{i=1}^{d}\Big(\tfrac{\sigma_{\mathbf{q},i}^{2}}{\sigma_{\mathbf{c},i}^{2}}+\tfrac{\sigma_{\mathbf{c},i}^{2}}{\sigma_{\mathbf{q},i}^{2}}-2\Big).$
Write each ratio using $\sigma_{\mathbf{q},i}^{2}=\tau+\delta_{\mathbf{q},i}$, etc., and expand:
$\frac{\sigma_{\mathbf{q},i}^{2}}{\sigma_{\mathbf{c},i}^{2}}=\frac{\tau+\delta_{\mathbf{q},i}}{\tau+\delta_{\mathbf{c},i}}=1+\frac{\delta_{\mathbf{q},i}-\delta_{\mathbf{c},i}}{\tau}-\frac{(\delta_{\mathbf{q},i}-\delta_{\mathbf{c},i})\delta_{\mathbf{c},i}}{\tau^{2}}+O(\varepsilon^{3}).$
Similarly for the reversed ratio. Summing and simplifying yields
$\tfrac{\sigma_{\mathbf{q},i}^{2}}{\sigma_{\mathbf{c},i}^{2}}+\tfrac{\sigma_{\mathbf{c},i}^{2}}{\sigma_{\mathbf{q},i}^{2}}-2=\frac{(\delta_{\mathbf{q},i}-\delta_{\mathbf{c},i})^{2}}{\tau^{2}}+O(\varepsilon^{3}).$
Hence
$V_{J}=\tfrac{1}{2\tau^{2}}\sum_{i=1}^{d}(\delta_{\mathbf{q},i}-\delta_{\mathbf{c},i})^{2}+O(\varepsilon^{3}).$
Note that each $\delta_{(\cdot),i}=O(\varepsilon)$, so $V_{J}=O(\varepsilon^{2})$.
Putting Mean And Variance Parts Together
Combining Eq. 21 and the expression for $V_{J}$ we get
$J=\tau^{-1}\cdot\tfrac{1}{2}\|\bm{\mu}_{\mathbf{q}}-\bm{\mu}_{\mathbf{c}}\|^{2}\;+\;O(\varepsilon)\;+\;O(\varepsilon^{2}).$ (22)
The dominant term is proportional to $\tfrac{1}{2}\|\bm{\mu}_{\mathbf{q}}-\bm{\mu}_{\mathbf{c}}\|^{2}$ with coefficient $1/\tau$; the variance-only contribution is $O(\varepsilon^{2})$.
Relation Between $J$ And 2(1s^)2(1-$\hat${s})
Recall from Eq. 15 that
$2(1-\hat{s})\;=\;\|\bm{\mu}_{\mathbf{q}}-\bm{\mu}_{\mathbf{c}}\|^{2}+\tfrac{\pi}{8}\sigma_{s}^{2}+O(\sigma_{s}^{4}).$
Comparing this with Eq. 22, we see that the mean-dependent parts of $J$ and 2(1s^)2(1-$\hat${s}) agree up to a constant scaling factor $1/\tau$. Therefore, defining the scalar scaling factor
$\alpha:=\tau^{-1},$
we have
$\alpha^{-1}J=\tfrac{1}{2}\|\bm{\mu}_{\mathbf{q}}-\bm{\mu}_{\mathbf{c}}\|^{2}+O(\varepsilon)+O(\varepsilon^{2}).$
Consequently
[…]
$=-\tfrac{1}{2}\|\bm{\mu}_{\mathbf{q}}-\bm{\mu}_{\mathbf{c}}\|^{2}+O(\varepsilon)+O(\varepsilon^{2}).$ (23)
Multiplying both sides by 1-1 and rearranging gives a relation of the form
$2(1-\hat{s})=\alpha^{-1}J+O(\varepsilon)+O(\varepsilon^{2}).$
Under Assumption 12, the $O(\varepsilon)$ terms refine to $O(\varepsilon^{2})$ as in the Wasserstein case (the same elementwise cancellation logic applies to the linear-in-$\delta$ components), yielding:
Theorem 2 (Affine approximation to Jeffreys divergence).
Under Assumptions 13, there exists a scalar $\alpha>0$ and a constant $C>0$ independent of $d$ such that for sufficiently small $\varepsilon$,
$\big|\,\alpha^{-1}J(q,c)-2(1-\hat{s})\,\big|\leq C\cdot\varepsilon^{2}.$
Equivalently,
$\hat{s}=1-\tfrac{1}{2}\alpha^{-1}J(q,c)+O(\varepsilon^{2}).$
A.3.5 Ranking Equivalance
Combining Theorem 1 (Wasserstein affine bound) and Theorem 2 (Jeffreys affine bound), we conclude that under the stated small-variance and common-scale assumptions the estimator $\hat{s}$ provides an affine, $O(\varepsilon^{2})$-accurate surrogate for both $W_{2}^{2}$ and (up to a constant scaling) the Jeffreys divergence; therefore, $\hat{s}$ induces the same ranking as these distributional distances up to $O(\varepsilon^{2})$ errors.
However, our bound relies on first-order Taylor expansions and variance dispersion assumptions, yielding $O(\varepsilon^{2})$ residuals. If these assumptions are violated (e.g., with highly concentrated variances or heterogeneous scales), the guarantee can degrade to $O(\varepsilon)$. Developing refined analyses that relax these conditions and provide tighter bounds remains an important direction for future work.
Appendix B Experimental Details and Additional Results
B.1 Preliminary Analysis of Embedding Subspace Alignment
Before ensemble the embeddings, a potential concern is whether embeddings from different models indeed lie in a sufficiently aligned subspace. We argue that in our setting this assumption is reasonable and empirically supported.
First, the embedding models considered in our experiments share (i) the same underlying architecture and (ii) largely overlapping training data. As established in prior work (Hewitt & Manning, 2019), such conditions encourage independently trained models to discover consistent linguistic structures—such as syntax, semantic relations, and contextual dependencies—and to encode them in geometrically comparable ways in their embedding spaces.
Second, our ensembling procedure applies $\ell_{2}$ normalization to embeddings before aggregation. This post-processing step has the effect of further mitigating potential subspace differences, effectively aligning embeddings to a common hypersphere and making their scales directly comparable (Timkey & Van Schijndel, 2021).
Finally, before ensembling, we empirically assessed the degree of subspace overlap by computing Canonical Correlation Analysis (CCA) (Hotelling, 1992) between models. We considered two model groups: (i) E5, BGE-zh, and MATRYOSHKA (Section 4.1), and (ii) BGE, E5, and GTE (Section 4.2). Since CCA quantifies the similarity and alignment between two sets of representations, we computed pairwise scores within each group of three models. For the first group, CCA was measured using the same dataset as in Section 4.1. For the second group, it was measured on Wikipedia dataset, which were used in the Retrieval task. As shown in Table 6, the results indicate that embeddings exhibit substantial overlap, supporting the view that they inhabit approximately aligned subspaces.
Table 6: Analysis of embedding subspace overlap using CCA (%). Results are reported for pairs in two different model groups.
Model Group CCA $\uparrow$ Model Group CCA $\uparrow$
(E5, BGE-zh) 99.9 % (BGE, E5) 99.9 %
(E5, MATRYOSHKA) 99.8 % (BGE, GTE) 99.4 %
(BGE-zh, MATRYOSHKA) 98.3 % (E5, GTE) 99.5 %
B.2 Comparison with Alternative Similarity Metrics
In this section, we compare our proposed uncertainty-aware similarity (UEC) against widely used probabilistic similarity metrics. We restrict our comparison to similarity functions that admit closed-form solutions between Gaussian distributions, thereby avoiding the computational overhead of Monte Carlo sampling. Specifically, we consider the KLD (Kullback & Leibler, 1951), 1-Wasserstein (Chhachhi & Teng, 2023), 2-Wasserstein (Gelbrich, 1990), and Closed-form sampled distance (CSD) (Chun, 2023).
KLD
Given two multivariate Gaussian distributions $\mathcal{N}_{1}=\mathcal{N}(\bm{\mu}_{1},\bm{\Sigma}_{1})$ and $\mathcal{N}_{2}=\mathcal{N}(\bm{\mu}_{2},\bm{\Sigma}_{2})$, the closed-form expression for KLD is:
$\mathrm{D}_{\mathrm{KL}}(\mathcal{N}_{1}\|\mathcal{N}_{2})=\frac{1}{2}\left[\mathrm{tr}(\bm{\Sigma}_{2}^{-1}\bm{\Sigma}_{1})+(\bm{\mu}_{2}-\bm{\mu}_{1})^{\top}\bm{\Sigma}_{2}^{-1}(\bm{\mu}_{2}-\bm{\mu}_{1})-d+\log\left(\frac{\det\bm{\Sigma}_{2}}{\det\bm{\Sigma}_{1}}\right)\right],$ (24)
where $d$ is the dimensionality of the distribution. While informative, this metric involves matrix inversion and log-determinants, making it computationally expensive for large-scale retrieval.
1-Wasserstein Distance
The 1-Wasserstein distance between two Gaussians $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ is given by:
$W_{1}(\mathcal{N}_{1},\mathcal{N}_{2})\approx\|\bm{\mu}_{1}-\bm{\mu}_{2}\|_{2}+\left|\mathrm{tr}(\sqrt{\bm{\Sigma}_{1}})-\mathrm{tr}(\sqrt{\bm{\Sigma}_{2}})\right|.$ (25)
Here, the distance accounts for both mean shift and scale (via trace of square-root covariance). Although it is more interpretable, it requires square-root matrix operations which are non-trivial.
2-Wasserstein Distance
The 2-Wasserstein distance between $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ is defined as:
$W_{2}^{2}(\mathcal{N}_{1},\mathcal{N}_{2})=\|\bm{\mu}_{1}-\bm{\mu}_{2}\|_{2}^{2}+\mathrm{tr}\left(\bm{\Sigma}_{1}+\bm{\Sigma}_{2}-2(\bm{\Sigma}_{2}^{1/2}\bm{\Sigma}_{1}\bm{\Sigma}_{2}^{1/2})^{1/2}\right),$ (26)
which also integrates both mean and covariance alignment. However, its computation is notably expensive due to the presence of matrix square roots and requires full-rank covariance matrices.
CSD
The Closed-form Sampled Distance (CSD), proposed in PCME++ (Chun, 2023). A key advantage of CSD is its tractable closed-form, given by:
$\mathrm{CSD}(\mathcal{N}_{1},\mathcal{N}_{2})=\|\bm{\mu}_{1}-\bm{\mu}_{2}\|_{2}^{2}+\|\bm{\Sigma}_{1}+\bm{\Sigma}_{2}\|_{1},$ (27)
where $\|\cdot\|_{1}$ denotes the element-wise $\ell_{1}$-norm. CSD captures both distributional shift and total uncertainty spread in a simple additive form, and is more computationally efficient than metrics involving matrix square roots or inverses.
We evaluate the similarity metrics under the MIRACL subset retrieval task described in Section 4.1. As shown in Table 7, our proposed UEC achieves the highest performance across all metrics—nDCG@10, Recall@10, and AUC@10—while maintaining the lowest runtime and linear complexity in both memory and time. In contrast, KLD and 2-Wasserstein, while theoretically rich, incur significantly higher computational costs due to matrix inversions and square-root operations, making them less suitable for scalable systems. Notably, UEC does not require any sampling and operates entirely in closed-form, striking an ideal balance between efficiency and uncertainty-awareness. These results underscore the practical advantages of our similarity formulation for real-world deployment.
Table 7: Comparison of similarity metrics under the MIRACL evaluation (Section 4.1). “Rel.” and “Comp.” denote relative runtime (normalized to UEC=1.00) and asymptotic complexity per similarity computation.
Similarity nDCG@10 Recall@10 AUC@10 Memory Complexity Time Rel. Complex. KLD 58.14 78.33 87.60 $\mathcal{O}(KD+KD^{2})$ 1.48 $\mathcal{O}(KD^{2})$ 1‑Wasserstein 57.87 88.21 79.92 $\mathcal{O}(KD)$ 1.21 $\mathcal{O}(KD^{1.5})$ 2‑Wasserstein 57.71 88.48 81.05 $\mathcal{O}(KD+KD^{2})$ 1.60 $\mathcal{O}(KD^{2})$ CSD 58.23 79.10 86.92 $\mathcal{O}(KD)$ 0.88 $\mathcal{O}(KD)$ UEC (Ours) 59.65 80.07 91.04 $\mathcal{O}(KD)$ 1.00 $\mathcal{O}(KD)$
B.3 Experimental Details: MIRACL Subset
B.3.1 Fitting Laplace Approximation
To quantify uncertainty for each embedding model, we apply a diagonal Laplace approximation (LA) to the final layer using a small, labeled dataset. In this experiment, we fit the approximation using the MIRACL Hard Negative (Zhang et al., 2023) Subset. For E5, we use 50 labeled examples each from English and Russian; for BGE-zh, we use 50 examples each from English and Chinese; and for MATRYOSHKA, we use 50 examples each from English and Arabic.
Each query-passage pair is assigned a binary label indicating whether the passage is relevant (1) or not (0) to the query. These binary-labeled pairs are then used to train a logistic classifier with a cross-entropy objective.
To ensure computational scalability, we adopt a diagonal Gaussian approximation. We fit the LA using the laplace-redux library [^2] https://aleximmer.com/Laplace/. The posterior mean corresponds to the MAP estimate of the final layer’s weights, and the posterior variance is derived from the inverse of the diagonal Hessian. This provides a lightweight probabilistic embedding representation without requiring backpropagation or full model fine-tuning.
B.3.2 Dataset
For the MIRACL subset evaluation, we constructed a test set by sampling 25 queries each from four languages (English, Russian, Chinese, and Arabic) from the multilingual MIRACL Hard Negative dataset (Zhang et al., 2023). This selection aimed to cover diverse language families and scripts, ensuring a balanced multilingual evaluation.
B.3.3 Metrics
We employed three commonly used retrieval metrics: nDCG@10, Recall@10, and AUC@10.
  • nDCG@10 (Normalized Discounted Cumulative Gain) measures the ranking quality based on the positions of relevant passages among the top 10 retrieved items. It assigns higher weights to relevant items appearing earlier in the list, penalizing misranked relevant results logarithmically.
  • Recall@10 evaluates whether at least one relevant passage is retrieved within the top 10. It is computed as the fraction of queries for which a relevant document appears among the top-10 predictions.
  • AUC@10 (normalized Area Under the Curve) measures how well a model’s confidence scores align with its actual retrieval performance when abstention is allowed. Specifically, the model is allowed to abstain on queries with low confidence scores, and AUC@10 tracks how the average performance improves as more uncertain queries are excluded. We compute AUC@10 by plotting nDCG@10 as a function of the abstention rate—i.e., the proportion of low-confidence queries discarded—and measuring the area under this curve. To enable fair comparison across models, we normalize this area between two bounds: (1) a baseline curve where abstention does not improve performance, and (2) an oracle curve where the worst-performing queries are perfectly abstained first. The normalized AUC is then computed as:
    $\text{nAUC}=\frac{\text{AUC}_{\text{Actual}}-\text{AUC}_{\text{Baseline}}}{\text{AUC}_{\text{Oracle}}-\text{AUC}_{\text{Baseline}}}.$
    Intuitively, higher nAUC values indicate that the model’s confidence scores are well-calibrated: as the model abstains on low-confidence queries, its average performance improves meaningfully. Conversely, a negative nAUC implies that the model’s confidence scores are misleading—i.e., low-confidence queries were actually high-performing, or high-confidence queries were often wrong.
B.3.4 Hyperparameters
We search hyperparameters for weight ensemble and UEC (Ours). In UEC, we introduce a temperature parameter to sharpen the ensemble coefficients, and additionally incorporate a hyperparameter $\beta$ to scale the influence of the variance $\sigma_{s}^{2}$ in the uncertainty-aware similarity estimation.
  • Weight Ensemble: We performed a grid search over weight combinations such as $[0.1,0.1,0.8]$, $[0.1,0.2,0.7]$, $\ldots$, $[0.8,0.1,0.1]$, ensuring the weights sum up to 1 across the three models.
  • UEC (Ours): We fixed the temperature as $1.5$ for all experiments. We search $\beta$ in range of $[0.0001,0.001,0.01,0.1]$.
B.3.5 Auxiliary Results
Table 8: Performance on MIRACL Subset across models and combination methods. The best and second-best results per row are bolded and underlined. UEC shows the highest performance on average across all metrics.
Language / Metric Individual Models Ensemble UEC
E5 BGE-zh MATRYOSHKA Uniform Weighted
English
   nDCG@10 80.27 67.05 40.02 69.82 76.30 78.29
   Recall@10 97.19 90.50 54.83 92.66 97.80 97.98
   AUC@10 37.37 -28.31 -86.74 -24.38 6.71 86.72
Russian
   nDCG@10 39.12 12.25 5.96 11.65 21.78 26.61
   Recall@10 48.98 15.98 7.21 15.93 30.77 41.38
   AUC@10 60.18 38.22 -46.00 17.14 39.65 91.49
Chinese
   nDCG@10 17.74 75.16 0.73 39.17 35.26 60.68
   Recall@10 28.65 97.45 2.00 57.02 54.60 91.04
   AUC@10 60.99 -53.35 27.89 13.65 38.01 92.14
Arabic
   nDCG@10 13.36 0.64 71.47 66.47 61.97 73.18
   Recall@10 17.87 1.33 92.55 85.01 79.43 89.88
   AUC@10 67.45 42.06 85.74 9.62 39.12 93.82
Avg. nDCG@10 37.62 38.77 29.55 46.78 48.83 59.65
Avg. Recall@10 48.17 51.32 39.15 62.66 65.65 80.07
Avg. AUC@10 56.50 -0.34 -4.78 4.01 30.87 91.04
Table 8 shows detailed results for individual models and combination methods. UEC achieves the highest average performance across all metrics, with particularly strong gains in AUC.
B.4 Experimental Details: MMTEB
B.4.1 Fitting Laplace Approximation
To estimate uncertainty for each embedding model, we fit a diagonal Laplace Approximation (LA) to the last layer of each model using a small subset of labeled data. We choose datasets that are commonly used in the pre-training or fine-tuning of the considered embedding models to ensure consistency between model behavior and fitting data.
Specifically, we use the MS MARCO (Nguyen et al., 2016) and SNLI (Bowman et al., 2015) datasets for fitting. From MS MARCO (consisting of roughly 80K samples), we randomly select 3,983 query-passage pairs. From SNLI (550K examples), we use 3,775 sentence pairs. Each triplet consists of a query (or premise), a positive passage (or hypothesis), and a randomly sampled negative passage. We assign binary labels indicating whether a passage is relevant (1) or not (0) with respect to the query. These binary-labeled pairs are used to fit a logistic classifier using cross-entropy loss.
We adopt the diagonal Gaussian approximation to ensure computational efficiency, and perform LA fitting using the laplace-redux library. The mean corresponds to the MAP estimate of the final layer weights, and the (inverse) diagonal Hessian is used to compute the posterior variance. This yields a probabilistic representation of each embedding without requiring re-training or backpropagation through the base model.
B.4.2 Dataset
For broader multilingual evaluation, we conducted experiments on MMTEB (Muennighoff et al., 2022; Enevoldsen et al., 2025), leveraging the official benchmark codebase available at https://github.com/embeddings-benchmark/mteb. We mainly selected a representative subset of tasks from the MMTEB benchmark, including five tasks from Classification, five from Retrieval, and ten from STS. The datasets for each task were chosen to span multiple domains within the same task type to assess domain-robust performance.
B.4.3 Metrics
In MMTEB evaluations, we used the main score recommended by the benchmark as the primary evaluation metric for each task. Additionally, we reported commonly adopted metrics for each task type. For Retrieval, we reported nDCG@10, Recall@100, and AUC@10. For classification, we reported Accuracy, F1 score, and AUROC. The F1 score is the harmonic mean of precision and Recall, providing a balanced measure of classification performance, especially under class imbalance. AUROC, short for the area under the receiver operating characteristic curve, measures a model’s ability to distinguish between classes across varying decision thresholds. It is particularly suited for evaluating models under uncertainty, as it considers both sensitivity and specificity to assess confidence calibration. For STS tasks, we used Spearman correlation, the main metric in MMTEB, which measures the rank correlation between predicted similarity scores and human-annotated ground truth.
B.4.4 Hyperparameters
  • Weight Merging: We used grid search over weight combinations including $[0.1,0.1,0.8]$, $[0.1,0.2,0.7]$, $\ldots$, $[0.8,0.1,0.1]$, ensuring the weights sum to 1.
  • Task Arithmetic: We implemented task arithmetic following Ilharco et al. (2022), where the merged embedding is computed by linearly combining base model embeddings with the direction vector between two other models. Formally, given three models $A,B,C$, the merged embedding is:
    $\text{Embed}_{\text{merged}}=\text{Embed}_{A}+\alpha\cdot(\text{Embed}_{B}-\text{Embed}_{C}),$
    where $\alpha$ is a hyperparameter controlling the direction scaling. We performed grid search over $\alpha\in\{0.0001,0.001,0.01,0.1,1.0\}$.
  • Weight Ensemble: We used grid search over weight combinations including $[0.1,0.1,0.8]$, $[0.1,0.2,0.7]$, $\ldots$, $[0.8,0.1,0.1]$, ensuring the weights sum to 1.
  • UEC (Ours): The temperature was fixed at $1.8$ throughout the experiments. We search $\beta$, scaling parameter for $\sigma_{s}^{2}$, in range of $[0.0001,0.001,0.01,0.1]$.
B.4.5 Additional Tasks
In addition to the three representative tasks (Retrieval, Classification, STS), we also conduct experiments on other MMTEB tasks, including Bitext Mining, Clustering, and Reranking. For each of these tasks, we randomly select a subset of datasets and observe that the proposed UEC consistently achieves strong performance.
Bitext Mining
Table 9 summarizes performance on multilingual bitext mining across six representative datasets from MMTEB: Bornholm (Derczynski & Kjeldsen, 2019), Flores (Goyal et al., 2022), IN22Gen (Gala et al., 2023), IndicGenBench (Singh et al., 2024), NTREXB (Federmann et al., 2022), NorwegianCourts (Tiedemann & Thottingal, 2020). For evaluation, we report F1 scores following MMTEB’s standard setup.
Table 9: Comparison of Bitext Mining performance. The best and second-best results per dataset are bolded and underlined. UEC achieves the hightest average performance overall.
Individual Models Model Merging Ensemble UEC Dataset BGE E5 GTE GTE-MB Uniform Weighted TaskArith Uniform Weighted BornholmBitextMining 27.60 38.49 32.16 27.64 33.72 37.21 37.03 34.50 36.99 38.82 FloresBitextMining 13.03 19.42 13.66 17.48 15.93 18.21 19.38 15.89 18.38 19.55 IN22GenBitextMining 4.54 8.85 3.52 5.88 5.94 8.13 8.26 6.41 8.52 8.60 IndicGenBenchFloresBitextMining 3.72 6.91 3.47 4.50 4.82 6.55 4.91 4.99 6.30 6.66 NTREXBitextMining 18.78 27.68 19.77 24.54 23.08 26.34 25.10 22.51 26.01 27.72 NorwegianCourtsBitextMining 90.67 92.91 92.03 91.45 91.28 92.04 91.87 91.52 92.62 92.96 Avg. F1 26.39 32.37 27.44 28.58 29.13 31.41 31.09 29.03 31.47 32.39
Our proposed UEC method achieves the best average performance, outperforming both individual models and all ensemble or merging strategies. Notably, UEC obtains the top result on four out of six datasets, demonstrating its ability to robustly fuse knowledge from diverse pretrained models under high uncertainty.
Clustering
We measure the performance on multilingual clustering across eight representative datasets from MMTEB: ArXivHierarchical, BigPatent (Sharma et al., 2019), Biorxiv, MasakhaNEWS (Adelani et al., 2023b), SIB200 (Adelani et al., 2023a), StackExchange (Geigle et al., 2021), WikiClustering.
The V-Measure is an external clustering evaluation criterion defined as the harmonic mean of homogeneity and completeness (Rosenberg & Hirschberg, 2007). Homogeneity measures whether each cluster contains only members of a single class, while completeness measures whether all members of a given class are assigned to the same cluster. V-Measure ranges from 0 (no alignment) to 100 (perfect clustering), and is widely used in multilingual and scientific document clustering benchmarks such as MMTEB.
Table 10: Comparison of Clustering performance. The best and second-best results per dataset are bolded and underlined. UEC achieves the hightest average performance overall.
Individual Models Model Merging Ensemble UEC Dataset BGE E5 GTE GTE-MB Uniform Weighted TaskArith Uniform Weighted ArXivHierarchicalClusteringP2P 59.75 58.05 59.85 59.50 59.38 60.14 59.56 60.29 58.42 61.43 ArXivHierarchicalClusteringS2S 58.05 54.91 57.67 58.16 58.95 58.63 57.08 58.87 58.35 59.32 BigPatentClustering.v2 31.03 30.69 29.30 30.28 31.36 33.92 31.08 31.95 35.41 34.87 BiorxivClusteringP2P.v2 41.65 38.01 40.58 41.91 40.98 40.86 40.56 41.37 40.68 43.06 MasakhaNEWSClusteringS2S 40.47 41.11 39.41 37.32 41.08 42.36 41.37 40.21 42.63 43.07 SIB200ClusteringS2S 9.38 10.04 9.50 11.40 10.87 11.13 11.36 11.13 11.66 12.25 StackExchangeClustering.v2 59.08 55.99 59.26 61.31 60.18 61.78 61.12 60.54 61.44 62.06 WikiClusteringP2P.v2 24.93 25.02 24.87 25.47 25.16 25.74 25.30 25.55 25.61 26.23 Avg. V-Measure 40.54 39.23 40.06 40.67 40.99 41.82 40.93 41.24 41.77 42.79
Table 10 summarizes performance on clustering tasks. UEC achieves the best average performance, outperforming all individual models and ensemble baselines. Notably, it ranks first on 6 out of 8 datasets. In short, UEC consistently improves by leveraging uncertainty-calibrated similarity and adaptive ensemble weighting, enabling more accurate cluster assignment under distributional shift and model disagreement.
Reranking
Table 11 presents the results of reranking experiments on five benchmark datasets: Alloprof (Lefebvre-Brossard et al., 2023), T2 (Xie et al., 2023), VoyageMMarco (Clavié, 2023), WebLINXCandidates (Lù et al., 2024), WikipediaReranking. We evaluate performance using Mean Average Precision (MAP), a standard retrieval metric that reflects both the precision and the ranking quality of relevant items. MAP is computed as the mean of the average precision scores over all queries, where average precision measures how well relevant documents are ranked near the top. Higher MAP values indicate that relevant passages are not only retrieved but also correctly prioritized.
Table 11: Comparison of Reranking performance. The best and second-best results per dataset are bolded and underlined. UEC achieves the hightest average performance overall.
Individual Models Model Merging Ensemble UEC Dataset BGE E5 GTE GTE-MB Uniform Weighted TaskArith Uniform Weighted AlloprofReranking 62.30 64.82 66.54 68.72 65.65 66.04 65.86 65.65 67.65 69.14 T2Reranking 60.93 60.55 61.20 65.24 61.83 62.98 62.77 61.39 63.34 65.08 VoyageMMarcoReranking 30.98 33.03 31.37 37.41 32.08 33.13 31.93 31.76 33.47 36.89 WebLINXCandidatesReranking 13.72 10.84 13.51 18.02 13.88 14.73 14.32 13.73 15.49 17.65 WikipediaRerankingMultilingual 72.65 76.83 73.78 74.51 75.83 76.32 75.47 75.21 76.74 77.87 Avg. MAP 48.11 49.21 49.28 52.78 49.85 50.64 50.07 49.55 51.54 53.33
The proposed UEC method achieves the highest overall performance with an average MAP, outperforming all other methods. It ranks first on three datasets and second on the remaining two. These results demonstrate UEC’s effectiveness in capturing uncertainty and enhancing similarity estimation for high-precision reranking.
Appendix C Statement on the Use of Large Language Models
In this study, we use LLMs in a limited, supporting capacity. The usage was restricted to polish writing, specifically for grammatical corrections, sentence refinement, and improving overall writing consistency, and experimental codes.